### v1

# Patterns of Patterns III

Abstract.
Building on earlier installments, this paper re-examines the PLACARD pattern. We report on a series of workshops where PLACARD was used to scaffold collaborative reflection, speculative inquiry, and stimulate design pattern generation. These accounts are enriched by a comparison case: virtual workshops carried out with simple AI-based chatbots. We discuss limitations and lessons learned from both the human and multi-agent settings. We conclude by outlining a future development strategy at the intersection of AI agents, design patterns, and institutional governance.
Design Patterns, Pattern Languages, Action Reviews, Futures Studies, Causal Layered Analysis, Free Software, Peeragogy, Artificial Intelligence, Anticipation, Doughnut Economics
copyright: rightsretainedjournalyear: 2021doi: 10.XXXX/XXXXXXX.XXXXXXXconference: PLoP’25: Pattern Languages of Programs 2025; October 12–15, 2025; HILLSIDEbooktitle: PLoP’25, OCTOBER 12–15, HILLSIDE. Copyright 2025 is held by the author(s)price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Social and professional topicsccs: Software and its engineering Designing softwareccs: Software and its engineering Open source modelccs: Applied computing Operations researchccs: Computing methodologies Modeling and simulation
1. Introduction
In an era of increasing global interconnectivity, effective collaboration across cultural, geographical, technical, and disciplinary boundaries is essential. Traditional methods often falter when addressing the intricate layers of context inherent in distributed collaboration. To address these concerns, we have been exploring a re-theorisation of design patterns using the tools of social intelligence. First introduced in our paper “Patterns of Patterns: A Methodological Reflection on the Future of Design Pattern Methods” (PoP), the PLACARD pattern was tested in five U.S. and U.K. workshops as detailed in the sequel “Patterns of Patterns II” (PoPII) (Corneli et al., 2021; Corneli et al., 2023). To support that work, PLACARD was elaborated into a design pattern language. It has since been adapted to additional localized settings, engaging diverse participants in hybrid and in-person formats. These sessions leveraged PLACARD to bridge perspectives and envision future modes of collaboration in rapid prototyping sessions. We have also been interested in how PLACARD can be used at a slower pace, and in distributed contexts—for example, to scaffold learning in a long-running seminar series—or even more widely, across different spheres of engagement that we have described in this series of papers. In this paper, we establish another use case, by using PLACARD patterns with virtual artificial intelligence (AI) agents. We show that design pattern thinking is well-suited to use in developing epistemic AI, by which we understand artificial intelligence systems that model and manage uncertainty, provide explanations, engage in meta-cognition, and facilitate human-AI collaboration. By working between human and computer contexts, we are able to improve our understanding of the underlying patterns of knowledge acquisition, representation, and utilization. Generative AI tools contributed in other ways to this paper, by helping us make the first draft based on their analysis of the two initial papers in this series and our three summary reports. We then took the baton and revised the writing over many subsequent versions to arrive at the cohesive work you are now reading.
2. Patterns in Use
2.1. Background
In its basic design pattern format, PLACARD combines three complementary approaches for collective learning and design.
PAR (Project Action Review):
– A reflective process assessing intentions, actions, and outcomes to learn from experience.
CLA (Causal Layered Analysis):
– A futures studies technique analyzing issues across multiple layers—the familiar litany of well-known problems, the systems that underpin them, along with the prevailing worldviews, and deeper myths that constitute the root causes of these problems, along with the possibility of developing alternative perspectives.
DPL (Design Pattern Language):
– A framework for capturing and sharing reusable solutions in a structured Context-Problem-Solution format (or similar), further structured by links between such elements.
By combining these approaches, PLACARD enables teams to cycle through reflection, deep analysis, and pattern creation. However, this is not necessarily a serial process or a fixed cycle. For instance, the generation, use, and analysis patterns may be ongoing. In our work a PAR is typically conducted after a period of work, to reflect on progress made, however it also serves as an intention-setting exercises might happen before the next phase begins. The nominal function of a CLA is to analyze a situation, but just as often they are used to describe how that situation can change.
The three papers in the series to date can be brought into schematic alignment with these methods.
(PoP $\approx$ CLA) We used CLA to think about the pattern literature as a whole and to retheorise its workings.
(PoPII $\approx$ DPL) We elaborated PLACARD into a pattern language.
(PoPIII $\approx$ PAR) We take a deep look at how these patterns have been used in practice, and take stock of potential next steps in the overall trajectory of research.
Thinking about patterns in pattern-theoretic terms, we uncovered PLACARD’s modularity. Methods that achieve the same function as one of its components can be swapped in and out, without undermining the integrity of the whole. For example, PAR could be replaced with the US Army’s After Action Review (US Army, 1993)—which inspired its basic format in the first place—or by the Feedback Wheel devised by Janet Hurley (Real, 2008). CLA could be replaced by Triple-Loop Learning (Tosey et al., 2012) or Ecological Systems Theory (Bronfenbrenner, 2000). DPL could be replaced by Elinor Ostrom’s institutions (Ostrom, 1990), or by a catalogue of literary tropes, and so on. Each such selection comes with its own ideological orientation, and typical domain of application.
2.2. Implementation
We implemented PLACARD in multiple settings, tailoring our approach to each context:
AsianPLoP 2024 (Fujisawa, Japan):
Our “PLACARD Workshop” (Danoff et al., 2024) was a hybrid focus group with one onsite and two remote facilitators (via Jitsi). Participants conducted meta-analysis on the conference itself using CLA and DPL to uncover underlying challenges, identify linguistic barriers and explore growth opportunities. One challenge discussed was how to be more inclusive of countries across the Asian continent. The conference organizers took a positive step by holding the 2025 event in Taiwan, but we cannot state that location decision was influenced by our Focus Group. In 2024, onsite attendees helped overcome barriers by clarifying for confused participants in Japanese. This was the only event in this paper conducted in multiple languages. A growth opportunity discussed that learning and networking does not need to stop when the conference itself ends. It can continue afterwards as attendees cross-pollinate pattern design collectively. The PAR was then applied to close the Focus Group and capture the results. During the evening conference recap ceremony there was one summary delivered in English by a paper author and another in Japanese by an attendee.
EuroPLoP 2024 (Irsee, Germany):
A hybrid session titled “AI Future Envisioning with PLACARD” (Tedeschi et al., 2024) with one onsite and three remote facilitators. Participants employed the Transfeminist Tech Oracle card game as a creative prompt, then used CLA and PAR in tandem to explore speculative AI futures. Due to time constraints and abstraction load, some participants expressed confusion about the method structure, prompting later pattern refinement. A post-session PAR survey was used for reflection after the event.
PLoP 2024 (Washington, USA):
An in-person “Imagination Run Wild” workshop, with three onsite facilitators. Participants were invited to “Envision the Future of Collaboration” (Ayloo et al., 2024), using the same card game and CLA. Another post-session PAR survey was used for reflection after the event.
Experts Among Us 2025 (Keene, New Hampshire, USA):
A stand-alone meeting in-person that gave an opportunity to play the same card game with one facilitator. In advance we practiced the card game and the facilitator then preselected the cards for participants. Attendees were not part of the patterns community.  This session actually facilitated a group discussion and demonstrated collaboration. We learned that we had to explain the rules and definitions prior to playing the game.
A common challenge for all of these workshops is rapid skill acquisition. We plan to continue this workshop series, for instance with a EuroPLoP 2025 workshop session. As mentioned above, this is a good time to take stock of what we’ve learned so far.
2.3. Reflexive Analysis Using the PLACARD Pattern Catalogue
To analyze the workshops beyond surface-level outcomes, we applied a reflexive methodology grounded in the PLACARD pattern catalogue, presented in PoPII. This catalogue is structured into three interlinked categories:
A. Identifying Themes (“sensory”) B. Organizing Structure (“cognitive”) C. Making It Actionable (“motor”)
For each workshop, in this paper we use the A/B/C structure to retrospectively assess:
  • Which patterns were implicitly or explicitly activated,
  • Where breakdowns in understanding or facilitation could be explained by missing or misapplied patterns, and
  • Where new patterns could be introduced to resolve to unanticipated tensions.
A brief summary is given in Section 3 of this paper. This reflexive approach helped us make sense of divergent workshop dynamics, and the possibility for further patterns that could help make the workshops more rewarding in general.  For example, Name the Frame was identified in hindsight as something that could have improved onboarding. Likewise, Harvest Before You Close was seen as a missing step in several sessions, and this pattern would address missed opportunities for extemporaneous pattern generation.
Rather than enforcing a rigid rubric, the catalogue functioned as a scaffold for learning from variation. It allowed us to compare facilitation dynamics, participant responses, and methodological adaptations across sites—and, crucially, to refine the catalogue itself through lived experience.
2.4. A Note on Notation
As above, we will use all-caps (like PLACARD) for patterns introduced in other sources, and small-caps (like Name the Frame) for patterns introduced in this paper, details of which are in the Appendix.
2.5. Empirical and Design-Based Development of New Patterns
Across the 2024 and 2025 events, new patterns did not emerge fully formed. Indeed, only one new pattern surfaced and named explicitly during the workshops:
  • Have Multiple Leaders Onsite – prompted by the hybrid setup at EuroPLoP, where limited physical facilitation led to friction. The pattern proposes distributed facilitation roles to support hybrid presence, reduce coordination load, and maintain session energy.
Other new proto patterns crystallized in our retrospective analysis. For example:
  • Name the Frame – (a further specification of CONTEXT SETTING from PoPII) this pattern emerged from repeated facilitator observations that participants struggled to orient themselves without a clear method overview.
  • Borrow the Situation – inferred from participant improvisation in choosing and reinterpreting prompts when predefined situations felt mismatched or abstract.
  • Start with Sense – (generalizing DÉRIVE COMIX from PoPII) builds on the observation that embodied play or metaphorical framing was often the key to unlocking participation
In sum, while few patterns emerged as polished artifacts within the workshop frame, the pattern language itself evolved through encounter. The workshops provided empirical grounding for evaluating utility, exposing limits, and highlighting latent needs. This iterative, design-based process parallels how pattern languages grow in living communities—not through top-down invention, but through grounded, reflective refinement.
2.6. Developing a comparison case: Working with patterns computationally
We noticed that it is possible to ask ChatGPT to run simple multi-agent “virtual” workshops. These can be compared with the in-person workshops. Below are two lightly edited example prompts that were used for setting up these simulations.
Prompt 1 Given to Generative AI Program I want you to create a multi-agent system with 15 different ChatGPT-based agents.  I want you to assign each of them the name of a distinct animal, and each of the agents will have some of the typical characteristics of the animal it is named for.  Then I want you to walk through the 30 “Peeragogical Exercises for a Seminar on Epistemic AI” [elided] in a virtual seminar.  The final exam is a dialogue between these animals, about the following prompts.  Please have as much discussion as you’d like among the agents before you give me their final team presentation answering the questions. What is epistemic AI? How does mathematics relate to this vision? What does ‘peeragogy’ mean, and how does it relate to this vision? What kind of multi-agent has created this answer? Why is this virtual seminar of historical importance?
Prompt 2 Given to Generative AI Program Remember how we did some simple multi-agent workshops before? Let’s simulate the architecture described the paper with some ChatGPT-based agents. This time, rather than basing them on animals, we are going to base the agents on a diverse group of Research Fellows in AI. Let’s only consider 10 such agents. Please include the “Ant” role from the workshop with animals, she will record what happens in which workshop session. Please also include the “Dog” role from the animal workshop, who will act as a primary reporter. The Fellows will be set the task of designing a series of workshops using the PLACARD method and associated patterns, and the outcome of the series should be a detailed plan of work for a research project on epistemic AI. You should do deep research to make sure that the discussions are plausible.
3. Patterns as Technical, Epistemic, and Reflective Infrastructure
3.1. Reframing Patterns as Sense-Making Scaffolds
Patterns are often described in instrumental terms: reusable solutions to recurring problems in context. As we look into how they work, we see there’s more to it. One aspect of this is the use-versus-mention distinction: clearly, a physical alcove is not the same as Christopher Alexander’s “ALCOVES” pattern (Alexander et al., 1977). And at both levels, there’s a degree of explicitness. When does an indentation become an alcove? When does a pattern become part of a language?
The Start with Sense pattern characterizes various ways of inviting participants to begin inquiry with embodied, perceptual, or intuitive grounding. This creates conditions under which problems and possibilities become intelligible.
Across the A/B/C catalogue, many patterns introduced earlier and from the Appendix operate similarly:
  • Name the Frame helps surface the underlying structure of an activity or method.
  • Borrow the Situation allows participants to reinterpret or co-create the context of inquiry.
  • Meta-Reflection Loop supports collective understanding of what just happened and why.
These are cognitive footholds, ways of managing ambiguity and distributing attention. They help groups compose a shared reality, especially when dealing with complex or unfamiliar material.
From this perspective, the value of a pattern lies not just in its repeatability but in its sensitivity to situated meaning-making. A good pattern doesn’t just say “do this”; it says “look here, attend to this tension, notice this affordance.” It invites participants to see the world in terms of actionable differences.
Reframing patterns this way brings PLACARD into closer dialogue with traditions such as:
  • Phenomenology—emphasis on situated perception (Moran, 2000)
  • Participatory design—emphasis on co-articulation of problems (Spinuzzi, 2005)
  • Grounded theory—emphasis on emergent categorization (Glaser, 1992)
It also prepares the ground for computational application. For artificial agents, much as for humans, sense-making begins with activation of relevance—the ability to determine what matters in a given context. In PLACARD, patterns serve as activation scaffolds: semi-formal cues that guide what should be noticed, named, or negotiated. Each of its layers—A/B/C—corresponds to a distinct mode of learning. Crucially, design patterns—or, indeed, any substitute schema that plays the same role—achieve the function of linking these modes, enabling transitions between experience, inquiry, and abstraction.
We propose that patterns support at least three interdependent forms of learning:
  1. (1)
    Reflective Learning (Boyd and Fales, 1983): Patterns such as PROJECT ACTION REVIEW  and Meta-Reflection Loop support meaning-making after action. They help participants surface experiences, identify what shifted, and consider next steps. This mode is retrospective and grounded in collective memory and sense-making.
  2. (2)
    Design Learning (Burdick, 2012): Patterns like Borrow the Situation or Pattern the Play create scaffolds for reimagining problems, proposing alternatives, and testing ideas. Design learning is generative, situated in ambiguity. Patterns here serve as framing devices that guide creativity while constraining chaos.
  3. (3)
    Model Learning: Patterns like Name the Frame and Harvest Before You Close deploy abstraction—capturing structural insights from activity. This form of learning supports knowledge representation and the translation of local insights into portable conceptual tools (e.g., (Ariyaratne, 1977)).
Patterns enable movement between these modes. For example:
  • A reflection may generate a design insight that prompts reframing.
  • A design intervention may require formal modeling to share or extend.
  • A model may provoke a return to embodied or emotional grounding.
3.2. Summary of the evidence base from our in-person workshops
Table LABEL:workshop-summary summarizes the use of PoPII patterns in our workshops. Further details are the Appendix.
Table 1. Summary of patterns used in workshops and potential for improvement
Category/Patterns Used Well Needs Refinement
A./DÉRIVE COMIX, DO YOUR RESEARCH, CONTEXT SETTING, PAR Overrun — PAR provides shared cohesion, when used; our PAR-inspired survey has a different feel. Context-setting could also be more cohesive: Name the Frame. Explore more creative, embodied PAR.
B./GOING META, FACILITATOR and FUNCTIONAL ROLES Mixed — time + onboarding were limits. Introduce meta-language tools earlier and more clearly. Test whether embodied cues help people get to worldview/myth.
C./REINFUSE EXPERTISE Implicit — patterns used, not made. Stronger follow-through on actionable steps; Harvest Before You Close. Explicitly support pattern emergence from cards/music activities.
4. Design Patterns for Pattern-Competent Agents
The PLACARD methodology, as developed and refined across multiple participatory workshops, demonstrates the power of design patterns not only as tools for structuring action, but also as scaffolds for shared sense-making and reflection. Building on this foundation, we now consider a forward-looking design space: the development of pattern-competent artificial agents capable of engaging with design patterns as part of collaborative inquiry. These agents would be expected to:
  • Recognize situations where a pattern might apply,
  • Use flexiformal representations of pattern logic,
  • Operate as a collective, i.e., to form multi-agent systems (MAS)
  • Participate in reflection loops that evolve the infrastructure itself.
4.1. Recognizing Pattern-Relevant Situations
The capacity to recognize situations where a pattern might apply lies at the heart of pattern competence. This is not a trivial matching task. It requires the ability to perceive and interpret contextual cues, to identify when a recurring structure or problem-type is emerging, and to discern salience: which aspects of a situation matter, and why. In participatory settings, such recognition is informed by experience, tacit knowledge, and social interaction. When developing artificial agents, the challenge is to develop mechanisms—symbolic, statistical, or hybrid—that allow for similarly situated recognition. Here’s an example showing how this might work for today’s computer systems:
Example: A chatbot trained to handle customer service inquiries recognizes that a user is “frustrated,” even if the words don’t match prior examples exactly—but the tone, pacing, and word choice are similar to previously tagged “frustrated” interactions.
Mechanism: A statistical similarity model (e.g., embedding distance + fine-tuned classifier) matches this new situation to prior ones and adjusts the agent’s response accordingly (e.g., a more empathetic tone or an escalation offer).
Next-generation systems might deal with the same scenario in more flexible ways, for instance, not only on the basis of stimulus-response pairs, but more fluidly—for instance, understanding how it was that the customer became frustrated and initiating changes to the company policy accordingly. This more meaningful form of ‘pattern recognition’ requires readiness to see and make sense of structure, and to reframe experience in meaningful, actionable terms — rather than to set data aside as a block, or (which is much the same) to cope with it in a purely reactive way.
4.2. Flexiformal Representation of Pattern Logic
To support agent-based reasoning, patterns must be expressed in a format that is both structured and open-ended. Drawing on the modified pattern structure used in PLACARD—ContextIf … (But)… ThenBecause…—each clause can be understood as a layer of scoping and activation:
Context::
a broad frame (task domain, discourse setting) that activates relevant subsystems,
If condition::
a salient event or state within that frame,
But::
a tension or constraint that introduces conflict,
Then::
a candidate action or communicative move.
Because::
While logic and reasoning have been core topics in computer science from its outset (Sowa, 1999), further research is needed to determine what sorts of reasoning—clearly combining abductive, deductive, inductive, and analogical modes—and what types of logic—classical, fuzzy, modal, monotonic, non-monotonic—are used by humans to think about patterns.
For now, we refer to the overall approach as “flexiformal” (Kohlhase, 2012), because this leaves the door open to combining different formal systems in a suitably flexible fashion. Simply adding more patterns won’t be enough to implement human-like reasoning in pattern-capable systems. For one thing, the patterns need to be used to dynamically allocate cognitive resources.
4.3. Reflection Loops and the Evolution of Epistemic Infrastructure
To sustain meaningful pattern use, agents must also be able to reflect. This means embedding meta-level structures like the PAR into the environment itself. Reflection loops should support the capture of learning, the revision of pattern catalogues, and the emergence of new design insights.
In future architectures, agents and humans alike can annotate, revise, and remix patterns, and participate in the shared cultivation of ‘living’ knowledge systems, in which patterns are part of a generative infrastructure for collective intelligence. We’ve previously imagined patterns being used within a “Discovery Machine” (Tedeschi et al., 2024).
4.4. Rethinking the Pattern Catalogue
The PLACARD A/B/C pattern catalogue was created with human participants in mind. It emphasizes social affordances, embodied prompts, and facilitation moves. While some of these translate readily into multi-agent contexts—e.g., “Meta-Reflection Loop” or “Pattern the Play”—others rely on human sensory input or informal framing strategies that are not directly available to AI systems.
Thus, we propose a tripartite model of pattern types needed for pattern-competent agents, generalizing the layers we used previously.
1. Interactional Patterns (generalizing the sensory layer, e.g., PAR)
PAR structures collaborative reflection and coordination. Interactional patterns in general guide agents in playing roles, reflecting, and negotiating meaning in collaborative settings—essentially the micro-level behaviors that mirror PAR’s inquiry loop. Examples include:
  • Role Reveal: Adopt explicit roles in multi-agent dialogue.
  • Meta-Reflection Loop: Periodically engage in structured self-review.
  • Timebox the Core: Manage resource allocation and attention windows.
These patterns require some translation to make it understandable to the machine, e.g., encoding roles as functions in MAS architectures or structuring turn-taking rules.
2. Cognitive Patterns (expanding upon the existing cognitive layer, e.g., CLA)
CLA unpacks multiple layers of meaning—litany, system, worldview, myth. These cognitive patterns help agents reason across contexts and detect tensions, embedding depth and nuance in internal reasoning. They are interpretive structures, supporting internal reasoning and pattern use:
  • Pattern Activation: Given a context and goal, retrieve applicable patterns.
  • Tension Detection: Use heuristics to identify contradictions or bottlenecks.
  • Candidate Move Generation: Translate pattern suggestions into operational plans.
3. Infrastructural Patterns (generalizing the motor layer, e.g., DPL)
DPL captures reusable structures for action. These infrastructural patterns define persistent templates and system-wide protocols that support reuse, feedback, and evolution of the broader system design. For example these are 3 meta-patterns that could enable engagement with patterns:
  • Shared Memory Architecture: Persistent storage of pattern use and modification. Putting patterns where people can get them, like the catalogue, akin to the SCRAPBOOK from Patterns of Peeragogy (Corneli et al., 2015)
  • Feedback Rhythms: Protocols for scheduling reflection and pattern revision.
  • Pattern Diffusion: Mechanisms for propagating new or adapted patterns across agent collectives.
They govern the architecture of the agent system much like DPL governs the architecture of human practice.
4.5. Desiging Environments for pattern-competent MAS
In addition to broader contextual information, relevant multi-agent system (MAS) environments would include the social and cognitive dynamics of inquiry: coordination, negotiation, reflection. Such environments have the capacity to support multiple, interacting Baldwin effects—individual learning loops that influence each other over time as the agents and their interactions evolve (Red’ko et al., 2005). This distributed adaptation mirrors cultural evolution: strategies become shared, revised, and recontextualized. Rather than simply evaluating success in terms of task completion or efficiency, MAS environments invite questions about “how” solutions are created.
This reflection constitutes suggests a key amendment to PLACARD. Whereas PLACARD was described in classic psychological terms (Hilgard, 1980) — which are also taken over by roboticists (Cassimatis et al., 2004) — it does not specifically talk about the environment in which the agents it considers would be situated. In robot terms, for example, we would expect an undersea robot and a Mars rover to have very different functionality, even if they have the same theoretical underpinning. This way of thinking suggests that the PLACARD pattern language should be extended to be an “A/B/C/D catalogue”, where the fourth section contains environmental or developmental patterns. With this in mind, in addition to further detail on the candidate patterns for pattern-competent agents — as outlined above — the appendix also includes some patterns that would guide the design of their working environments.
5. Virtual Multi-Agent Workshops in the Spirit of PLACARD
To explore the idea of pattern-competent agents developed Section 4, we conducted a series of virtual workshops using fictional agents within ChatGPT.  Using the PLACARD method, those agents rotated through breakout sessions, engaged in pattern-guided inquiry, and made use of a curated and emergent pattern catalogue. It is worth emphasizing that the workshops were computer-aided design fictions (we have previously explored this way of working “by hand” (Danoff et al., 2019), and we’re aware that fictional works can have unexpected real-world ramifications, e.g., (Wikipedia contributors, 2025)). ChatGPT can simulate plausible dialogues, behaviors, and reflective loops based on design patterns, but it cannot build and run custom agent architectures. All ‘agents’ were textual personas that were simulated by one model. The agents were plausibly able to discuss the design patterns and employ them in light-weight design exercises. Below is an excerpt from one such workshop, in which several virtual agents begin collaboratively applying the PLACARD methodology to plan a research project on epistemic AI. The agents have somewhat stereotyped fictional biographies which to some extent shape their responses (e.g., “Carlos – Epistemology & Systems Philosophy”; “Mei – Participatory Design & Feminist STS”).
Excerpt: PLACARD Planning Dialogue on Epistemic AI Carlos: I propose we begin with CONTEXT SETTING. If we’re designing a research project on epistemic AI, we need a shared sense of what “epistemic” means in our context. Priya: Agreed. For me, it’s about how agents—human or artificial—come to know, not just calculate. So I’d emphasize design patterns that support interpretive, situated reasoning. Raj: Right. Maybe we could use Borrow the Situation—look at real-world cases of failure or success in knowledge systems. Misinformation management? Scientific discovery pipelines? Sofia: I think we need to Name the Frame explicitly. Are we researching tools for knowledge validation? Or institutional models for human-AI epistemic partnerships? Mei: Maybe both. What if we used Pattern the Play to design speculative workshop sessions? Each session could simulate a different epistemic tension—uncertainty, disagreement, novelty—and trial agent responses. Yuki: And we need some kind of Meta-Reflection Loop. A structure where, after each workshop, we reflect not just on what we learned, but how the patterns performed. This could feed into our developmental arc. Carlos: So the goal of the series isn’t just content generation—it’s to test, refine, and expand the pattern catalogue itself? Sofia: Exactly. And that aligns with the DO YOUR RESEARCH pattern: we’re generating grounded material and modeling the tools that structure knowledge formation. Raj: We’ll need a pattern-aware toolchain too—maybe a Pattern Knowledge Repository that evolves with each workshop. Otherwise we lose coherence. Priya: Let’s add a Timebox the Core phase to every session. This gives us just enough structure to avoid sprawling, without over-engineering. Mei: I’ll draft a plan using PLACARD as a scaffolding template. We can each take ownership of one layer—PAR, CLA, DPL—and develop sub-patterns and protocols that fit.
6. Discussion
6.1. Limitations and Lessons Learned
While the outcomes of our PLACARD-based workshops (both in-person and simulated) were promising, several limitations emerged that highlight areas for refinement—both in the method itself and in our implementation of it. Generally, these limitations correspond to lessons that could be applied in future work.
Limitation Lessons Learned
Cognitive demands of CLA: The abstraction and depth of Causal Layered Analysis (CLA) often overwhelmed participants, especially when scaffolding was limited. Provide lightweight onboarding tools—such as mini-patterns, worked examples, or scenario walkthroughs—to ease entry into layered analysis.
Short workshop windows (60–90 mins): Limited time forced trade-offs, with later phases (e.g., PAR) often rushed or skipped. Recognize Timebox the Core as a necessary pattern—but pair it with mechanisms for asynchronous reflection or follow-up synthesis.
Challenges in hybrid facilitation: Remote facilitators struggled to engage with group energy and activities; in-person facilitators were overburdened. Adopt and refine the pattern Have Multiple Leaders Onsite, and explore interactional choreography to balance facilitation across modalities.
Sequential framing constrained emergent inquiry: Participants sometimes forced their intuitive insights into retroactive justifications. Allow more modular, remixable process designs. Embed “Improvise to Test” and “Pattern the Play” to legitimize non-linear sense-making.
Deliberative bias limits affective and tacit knowledge: Structured methods sometimes crowded out emotional or disruptive forms of insight. Integrate methods or prompts that scaffold expressive, intuitive, and affective contributions—e.g., through metaphor, narrative, or art-based reflection.
Risk of premature formalism: Design patterns encouraged generalization at the expense of context specificity. Emphasize pattern annotations and situated stories. Use pattern revision cycles to adapt and diversify rather than standardize.
Logistical overhead of documentation: Without dedicated roles or tooling, PLACARD’s emphasis on shared surfaces was burdensome. Support tool development and role formalization (e.g., for pattern capture). Design light-touch alternatives for low-resource contexts.
Simulated agents were archetypal, not adaptive: The multi-agent models lacked real-time learning or memory; they played out scripted reasoning paths. Treat the current set of simulations as design probes rather than predictive models. Use them to explore design space and pattern viability, not behavioral realism.
Idealized dialogues lacked messiness: No true misalignment, conflict, or unresolved disagreement appeared in the simulations. Future work should inject structured friction—conflicting goals, values, or timing—to better stress-test interactional patterns.
No persistent identity or memory across sessions: Agents did not evolve across workshops, limiting insight into long-term learning. Incorporate persistent memory structures and episodic traces to study cumulative reflection, pattern evolution, and institutionalization.
Language model constraints shaped outcomes: Agent behaviors emerged from a coherence-biased generative system, not open-ended cognition. Simulations must be interpreted with epistemic humility. Their role is speculative, not definitive—ideal for surfacing design tensions, not modeling reality.
6.2. Related Work
Looking forward, the next phase of PLACARD’s evolution could be a platform for cultivating design-pattern-capable agents, communities, and learning environments. PLACARD should be understood not as a finished method but as a generative infrastructure: a living commons of epistemic tools, shaped by their user communities.  These insights can be combined with Elinor Ostrom’s Design Principles for Governing a Commons to serve as a lens for evaluating and extending PLACARD, beyond the “agent” level to the “environment” level. On this point, we can point to some work: Pitt et al. (2012) explore how institutions can self-organize by formalizing socio-economic principles that guide agent behavior within Multi-Agent Systems (MAS). Drawing inspiration from Elinor Ostrom’s work, they propose an axiomatic design approach that captures principles such as proportionality, accountability, and collective choice in a machine-interpretable form. Where Pitt et al. emphasize axioms as foundational building blocks, our PLACARD-based approach leverages design patterns as intermediate representations that bridge human practice and machine reasoning.  In a more recent paper, Mertzani and Pitt (2024) present a methodological framework for integrating power sensitivity into the design of cyber-physical systems by drawing on second- and third-order cybernetics. They propose pattern-based design methods that foreground the observer’s position, across different layes of embeddedness—with beneficial applications of technology ranging from “Saves Physical and Cognitive Resources” to “Solves Complex Problems” to “Deals with Social Constructions” to “Mitigates Existential Crises”.
7. Conclusion
Looking back over the limitations and lessons learned, it’s worth drawing out one top-level remork. ChatGPT is adept at text processing, and we have put it to use significantly in our work on this paper. It was able to generate and discuss design patterns that would describe pattern-capable agents. However, it was not able to actually use these patterns to build and run pattern-capable agents. We hope others will be interested in using LLMs to work with design patterns—but we also hope that such usage will serve as a landmark rather than a destination.
Including a computational use case within the paper suggested a key amendment to PLACARD: adding a fourth environment dimension to the pattern catalogue . Human-facing and Human-Computer Interface-facing patterns should eventually be extended similarly.
For now, to close this paper, we put forward the following CLA, which brings together the “human” and “AI” use cases and outlines a trajectory of future work. In the context of this paper, the CLA layers can aligned with the expanded view of PLACARD that we have discussed—encompassing sensory (litany), cognitive (system), motor (worldview), and environment (myth) dimensions.
Litany: Known Problems and Existing Solutions
These are the most visible and widely discussed elements of current AI discourse: the promise and perils of large language models (LLMs) and the early success of LLM-associated code-writing agents.
Language Model Substrate
These foundational models (e.g., GPT-4) provide generative capabilities, statistical reasoning, and natural language interfaces. However, they lack structure, grounding, and memory.
Code-Writing & Execution Agents
Many simple agents today (e.g., auto-scripting agents, plugin interfaces) use LLMs to write and execute code or perform isolated tasks. These are effective but brittle, with limited reasoning transparency or contextual learning.
Systemic Layer: Epistemic Infrastructure via Patterns
To move beyond isolated capabilities, we introduce pattern-based epistemic scaffolding for multi-agent collaboration. This makes the system reflexive, composable, and governed by design principles.
Pattern-Aware Dialogue Agents
Agents trained to recognize and apply design patterns (from the PLACARD catalogue) in their interactions. They exhibit situated competence and can engage in structured reasoning, scenario exploration, and inquiry.
Pattern-Reflective Meta-Agents
These agents do not just use patterns—they also monitor, revise, and propose new ones, participating in iterative design and reflective cycles. They enable epistemic learning at the system level, coordinating multi-agent behavior across contexts.
Worldview Layer: Institutional Design through MAS
Inspired by Elinor Ostrom’s principles for governing commons, we treat the MAS as a social-technological institution that evolves through interaction.
Multi-Agent Institution Designers
This layer operationalizes developmental patterns that support reflexivity, feedback, nested structures, and shared norm evolution across MAS environments. These agents co-design systems of engagement, monitor norm congruence, and resolve governance disputes—just as human institutions do. They instantiate a worldview in which intelligence is not just personal or cognitive, but systemic and collective.
Myth Layer: Human Institutions as Living Models
Beyond technical systems, our proposal invites reflection on how real-world institutions—academic, policy, scientific—can themselves become more adaptive, transparent, and pattern-literate through interaction with these tools.
Real-World Interface
This is the last layer because it is grounded in recursive reform. Humans build tools to simulate and improve human institutions—not to replace them, but to co-evolve. This layer anchors the entire architecture in civic legitimacy, ethical accountability, and reflective participation. Our claim is not merely that we can control the system, but that we can use it to help re-pattern the world we inhabit. By working with agents that are both pattern- and institution-capable, we gain assurances not only about avoiding “run-away AI” scenarios, but about avoiding other tragedies of the commons.
8. Acknowledgements
We wish to acknowledge discussions with Cris Viseu around our 2025 EuroPLoP workshop, which continues the line of work described in the paper.
This paper also benefited from extensive interaction with OpenAI’s ChatGPT. Its capacities for text generation, summarisation, pattern elicitation, and recursive prompting played a formative role in shaping both the structure and substance of this work. We also gratefully acknowledge the unnamed authors of the texts on which ChatGPT was trained. Responsibility for the interpretations, appropriations, and conclusions presented here remains entirely with the human authors.
APPENDIX
This appendix supports the analysis in the paper. It contains:
  1. (1)
    A pattern-based analysis of the 2024 workshops, using the patterns from PoPII, and exploring the extent to which these patterns were used (implicitly or explicitly). A summary appears in Table LABEL:workshop-summary in Section 3.2.
  2. (2)
    Potential new PLACARD patterns, mentioned in the paper.
  3. (3)
    Candidate Patterns for Pattern-competent agents.
  4. (4)
    Candidate Patterns for Pattern-competent Agent Environments.
Pattern analysis of the 2024 workshops
One immediate result was the identification of of a proto-pattern, “Have Multiple Leaders Onsite,” addressing challenges in hybrid-facilitated events:
Context: Interactive conference workshops with both in-person and remote participants.
Problem: A single onsite facilitator struggles to manage breakout groups, technology, and overall flow, leading to gaps in support and reduced engagement (especially for remote attendees).
Solution: Deploy multiple co-facilitators physically onsite to coordinate activities, handle logistics, and engage with participants. Onsite leaders can divide roles (e.g., one per breakout table or task), ensuring better guidance and responsiveness, while remote facilitators focus on online participants.
A. Identifying Themes
DÉRIVE COMIX
Application: Participants used cards (object, value, body/territory, situation) as interpretive prompts to spark imagination—mirroring urban dérive in a conceptual sense. Observation: The card game facilitated divergent thinking akin to exploring unfamiliar terrain.
CONTEXT SETTING
Application: Early slides and explanations introduced key concepts (e.g., design patterns, CLA). Opportunity: Participant feedback highlighted a need for clearer upfront orientation. Improved alignment with this pattern could increase accessibility.
DO YOUR RESEARCH
Application: Past workshop experiences were integrated into the design; earlier uses of CLA and PAR informed this session. Example: Reference to Anticipation 2019 and prior events illustrates “pre-research” informing present design.
B. Organizing Structure
GOING META
Application: The project itself exemplifies recursive methodology—using PLACARD to refine PLACARD. Highlight: Participants were prompted to reflect on and imagine both the methodology and the subject (AI futures).
PATTERN LANGUAGE COMPONENTS
Partial: The workshop design showcases layering (CLA + DPL + PAR) but participants struggled to parse the conceptual load quickly. Recommendation: Use lighter keywords (HOWEVER, BECAUSE, SPECIFICALLY) to scaffold understanding incrementally.
FACILITATOR ROLES / FUNCTIONAL ROLES
Application: Hybrid facilitation with distinct roles (onsite leader, remote supporters) models distributed facilitation. Opportunity: Could be made more explicit—e.g., naming “Linker,” “Reflector,” or “Wrinkler” roles to participants.
C. Making it Actionable
THE FUTURE BEGINS NOW
Application: The card game generated imagined futures that participants shared and developed. Gap: While creative ideas emerged, follow-through actions or small-scale pilots were not described. An actionable step could have enhanced continuity.
PROJECT ACTION REVIEW (PAR)
Application: Strong use of the PAR pattern at the end of the session to reflect, share learnings, and gather insights. Insight: Participants gained confidence and intended to reuse techniques—an indicator of success.
REINFUSE EXPERTISE
Application: Participants brought diverse professional and cultural perspectives. Expertise was implicitly reintegrated via the shared reflection on AI futures and governance. Suggestion: A more explicit arc—temporarily setting aside expertise to foster common ground—could enrich this pattern’s application.
Potential new PLACARD patterns
A. IDENTIFYING THEMES
Start with Sense
Context: You are facilitating a workshop intended to generate new ideas or shared understanding. If: You want to activate creativity and presence, But: participants arrive cold or disconnected from each other and the topic, Then: Begin with a sensory or embodied activity (e.g., rhythm, image-based storytelling, gesture) that creates shared energy and primes pattern recognition. Evidence base from related literature: Techniques such as ”story circles” (participatory arts), embodied methods in design ethnography, and rhythm-based facilitation in music education all affirm the value of grounding group attention through shared sensory engagement. These strategies mirror the function of PAR alternatives that aim to surface and align affective states as a precondition for inquiry.
Name the Frame
Context: You are introducing participants to a method that is unfamiliar or complex (like PLACARD). If: You want them to participate actively, But: the framing is opaque or abstract, Then: Explicitly describe the method, its purpose, and how today’s session fits into the broader picture. Evidence base from related literature: This mirrors practices from facilitation toolkits like Liberating Structures (e.g., ”Purpose-to-Practice”) and the role of explicit framing in adult learning theory, where naming the structure of engagement is key to building learner autonomy and motivation.
Borrow the Situation
Context: You are using prompts (cards, challenges, scenarios) to guide group inquiry. If: You offer a situation that feels flat or generic, But: participants want something meaningful to anchor their play, Then: Invite participants to select, remix, or define their own situation to explore. Evidence base from related literature: User-centered design emphasizes co-defining the design challenge. Futures methods such as the ”Futures Triangle” and participatory foresight often encourage the co-creation of situational prompts to enhance contextual fit and participant investment.
B. Organizing Structure
Pattern Priming
Context: You are working with participants who may be new to design patterns. If: You want people to identify and name patterns, But: the idea of a ”pattern” is still abstract, Then: Seed early examples or light prompts (”What did you do just now that worked?”) to build fluency. Evidence base from related literature: This aligns with double-loop learning (Argyris and Schön), where patterns of thought and action are surfaced through inquiry. It’s also reinforced in cognitive apprenticeship, which models strategies before expecting learners to apply them.
Timebox the Core
Context: You have a limited workshop time (e.g., 60–120 minutes). If: You want to run through all three PLACARD components, But: you risk running out of time or overwhelming participants, Then: Identify and focus on a minimal viable loop (e.g. Situation → Scenario → Reflection) and simplify supporting activities. Evidence base from related literature: Agile retrospectives and timeboxing practices in design sprints demonstrate how temporal constraints support clarity, focus, and iterative learning without requiring completion of all intended steps.
Role Reveal
Context: You’re running a hybrid or multi-facilitator session. If: Facilitation roles are unclear, But: smooth coordination is needed to support the flow, Then: Make the roles explicit (e.g., Linker, Reflector) and let participants know who does what. Evidence base from related literature: In distributed team management and collaborative learning, named roles (e.g., Scrum Master, Scribe) are shown to enhance group performance and task clarity. This also connects to facilitation protocols in online and hybrid pedagogies.
C. Making it Actionable
Harvest Before You Close
Context: You’re running an engaging session. If: Participants have surfaced great insights, But: waiting until the end to name patterns might mean running out of time, Then: Add a short mid-session or pre-wrap-up step to ask: “What’s something you did or saw that could be reused elsewhere?” Evidence base from related literature: After Action Reviews (military, management), Feedback Wheels (relationship therapy), and Agile retrospectives all highlight the value of capturing learning as it emerges—not merely as a concluding gesture.
Pattern the Play
Context: You’re using creative methods (cards, music, improv) as part of your workshop. If: These are treated as energizers or tangents, But: you also want them to inform the core inquiry, Then: Treat playful activities as valid epistemic inputs. Ask: “what design moves or values emerged here?” Evidence base from related literature: Design ethnography, improv for innovation, and serious games scholarship all provide grounding for the idea that nonverbal, embodied, and affective practices contribute directly to sense-making and design thinking.
Meta-Reflection Loop
Context: You’re using a structured method like PLACARD itself. If: You reflect only on the topic (e.g., music and AI), But: you find yourself neglecting the method, Then: Add a closing question or section to reflect on the method’s affordances, limits, and what to revise. Evidence base from related literature: This is consistent with reflective practice (Schön), meta-evaluation in participatory action research, and “method feedback” sessions used in critical design studies. It also draws from Buddhist and contemplative traditions of meta-awareness and iterative refinement (e.g., Dependent Origination).
Candidate Patterns for Pattern-competent agents
Pattern Activation
Context: An agent receives a new task or encounters a novel situation. If: The system has access to a pattern catalogue relevant to the domain, But: Not all patterns are applicable or salient, Then: Retrieve candidate patterns using contextual filters (e.g., task type, known tensions) to prime downstream reasoning.
Tension Detection
Context: An agent is assessing a situation, proposal, or evolving dialogue. If: Indicators suggest divergence, contradiction, or friction, But: These tensions remain unmarked and unresolved, Then: Flag the tension explicitly and propose an exploratory frame (e.g., apply CLA or consult conflict-resolution patterns) to surface underlying issues.
Candidate Move Generation
Context: An agent has identified a situation in which one or more patterns apply. If: The context supports multiple interpretations or strategies, But: There’s uncertainty about how to instantiate the pattern concretely, Then: Generate a set of plausible candidate actions based on pattern logic, rank or cluster them using salience heuristics, and propose next steps.
Shared Memory Architecture
Context: Multiple agents operate in a dynamic environment over time. If: Patterns are used, modified, or co-created through interaction, But: There is no mechanism for recording or retrieving this history, Then: Store pattern instances, adaptations, and usage contexts in a persistent shared memory, enabling learning and reuse.
Feedback Rhythms
Context: A system includes agents collaborating over long timeframes or iterations. If: Patterns and strategies become outdated or misaligned, But: There is no regular cadence for reviewing or revising them, Then: Schedule periodic feedback sessions or checkpoints where agents reflect on outcomes, revise patterns, and tune system-wide epistemic strategies.
Pattern Diffusion
Context: A new pattern proves effective within a local agent subgroup. If: Other agents face similar challenges or tasks, But: There’s no pathway for knowledge transfer or generalization, Then: Broadcast the pattern across the system with metadata (e.g., domain, confidence score, context) to allow others to trial, adapt, or contest it.
Candidate Patterns for Pattern-competent Agent Environments
Institutional Drift
Context: Repeated MAS sessions produce cumulative outcomes. If: Pattern use becomes routinized, But: Changing contexts or challenges are not reflected in pattern updates, Then: Trigger a meta-level PAR to recontextualize relevance. Because: Institutions require periodic re-grounding to avoid ossification.
Norm Seed
Context: A new MAS instance is initiated. If: There are no agreed-upon norms, But: Some agents possess relevant precedents or analogues, Then: Seed group interaction with foundational mini-patterns. Because: Initial framing influences long-term behavior and coherence.
Boundary Oscillation
Context: MAS agents engage across different domains or networks. If: Norms, roles, or goals diverge significantly, But: Agents have memory of shared pattern lineage, Then: Engage a boundary negotiation pattern to recalibrate alignment. Because: Distributed epistemic systems must periodically reconverge.
Reflection Scaffold Upgrade
Context: Existing reflection protocols underperform in capturing learning. If: Meta-cognition fails to yield new patterns, But: System logs or feedback signals suggest deeper insight is possible, Then: Introduce upgraded reflection scaffolds (e.g., layered PAR, narrative loops). Because: Developmental learning depends on scaffolding that evolves with system maturity

### v2

@@
-Patterns of Patterns III
+Patterns of Patterns III: Agentic Patterns
 Abstract.
-Building on earlier installments, this paper re-examines the PLACARD pattern.
-We report on a series of workshops where PLACARD was used to scaffold collaborative reflection, speculative inquiry, and stimulate design pattern generation.
-These accounts are enriched by a comparison case: virtual workshops carried out with simple AI-based chatbots.
-We discuss limitations and lessons learned from both the human and multi-agent settings.
-We conclude by outlining a future development strategy at the intersection of AI agents, design patterns, and institutional governance.
-Design Patterns, Pattern Languages, Action Reviews, Futures Studies, Causal Layered Analysis, Free Software, Peeragogy, Artificial Intelligence, Anticipation, Doughnut Economics
-copyright: rightsretainedjournalyear: 2021doi: 10.XXXX/XXXXXXX.XXXXXXXconference: PLoP’25: Pattern Languages of Programs 2025; October 12–15, 2025; HILLSIDEbooktitle: PLoP’25, OCTOBER 12–15, HILLSIDE.
+Through a series of workshops, we looked at ways to structure and scaffold group dialogue, and support the emergence of novel design patterns.
+We contrast these sessions—which we ran with other humans—with two “virtual workshops” which we simulated with ChatGPT.
+Limitations in both human and virtual settings are discussed, alongside lessons learned.
+We conclude by proposing a development trajectory that combines AI agents, pattern-based design, and institutional governance.
+Design Patterns, Pattern Languages, Artificial Intelligence
+copyright: rightsretainedjournalyear: 2025doi: 10.XXXX/XXXXXXX.XXXXXXXconference: PLoP’25: Pattern Languages of Programs 2025; October 12–15, 2025; HILLSIDEbooktitle: PLoP’25, OCTOBER 12–15, HILLSIDE.
 Copyright 2025 is held by the author(s)price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Social and professional topicsccs: Software and its engineering Designing softwareccs: Software and its engineering Open source modelccs: Applied computing Operations researchccs: Computing methodologies Modeling and simulation
 1. Introduction
-In an era of increasing global interconnectivity, effective collaboration across cultural, geographical, technical, and disciplinary boundaries is essential.
-Traditional methods often falter when addressing the intricate layers of context inherent in distributed collaboration.
-To address these concerns, we have been exploring a re-theorisation of design patterns using the tools of social intelligence.
-First introduced in our paper “Patterns of Patterns: A Methodological Reflection on the Future of Design Pattern Methods” (PoP), the PLACARD pattern was tested in five U.S.
+PLACARD is a design pattern for building and using design patterns (Corneli et al., 2021).
+The pattern was explored in five U.S.
 and U.K.
-workshops as detailed in the sequel “Patterns of Patterns II” (PoPII) (Corneli et al., 2021; Corneli et al., 2023).
-To support that work, PLACARD was elaborated into a design pattern language.
-It has since been adapted to additional localized settings, engaging diverse participants in hybrid and in-person formats.
-These sessions leveraged PLACARD to bridge perspectives and envision future modes of collaboration in rapid prototyping sessions.
-We have also been interested in how PLACARD can be used at a slower pace, and in distributed contexts—for example, to scaffold learning in a long-running seminar series—or even more widely, across different spheres of engagement that we have described in this series of papers.
-In this paper, we establish another use case, by using PLACARD patterns with virtual artificial intelligence (AI) agents.
-We show that design pattern thinking is well-suited to use in developing epistemic AI, by which we understand artificial intelligence systems that model and manage uncertainty, provide explanations, engage in meta-cognition, and facilitate human-AI collaboration.
-By working between human and computer contexts, we are able to improve our understanding of the underlying patterns of knowledge acquisition, representation, and utilization.
-Generative AI tools contributed in other ways to this paper, by helping us make the first draft based on their analysis of the two initial papers in this series and our three summary reports.
-We then took the baton and revised the writing over many subsequent versions to arrive at the cohesive work you are now reading.
-2. Patterns in Use
-2.1. Background
-In its basic design pattern format, PLACARD combines three complementary approaches for collective learning and design.
-PAR (Project Action Review):
-– A reflective process assessing intentions, actions, and outcomes to learn from experience.
-CLA (Causal Layered Analysis):
-– A futures studies technique analyzing issues across multiple layers—the familiar litany of well-known problems, the systems that underpin them, along with the prevailing worldviews, and deeper myths that constitute the root causes of these problems, along with the possibility of developing alternative perspectives.
-DPL (Design Pattern Language):
-– A framework for capturing and sharing reusable solutions in a structured Context-Problem-Solution format (or similar), further structured by links between such elements.
-By combining these approaches, PLACARD enables teams to cycle through reflection, deep analysis, and pattern creation.
-However, this is not necessarily a serial process or a fixed cycle.
-For instance, the generation, use, and analysis patterns may be ongoing.
-In our work a PAR is typically conducted after a period of work, to reflect on progress made, however it also serves as an intention-setting exercises might happen before the next phase begins.
-The nominal function of a CLA is to analyze a situation, but just as often they are used to describe how that situation can change.
-The three papers in the series to date can be brought into schematic alignment with these methods.
-(PoP $\approx$ CLA) We used CLA to think about the pattern literature as a whole and to retheorise its workings.
-(PoPII $\approx$ DPL) We elaborated PLACARD into a pattern language.
-(PoPIII $\approx$ PAR) We take a deep look at how these patterns have been used in practice, and take stock of potential next steps in the overall trajectory of research.
-Thinking about patterns in pattern-theoretic terms, we uncovered PLACARD’s modularity.
-Methods that achieve the same function as one of its components can be swapped in and out, without undermining the integrity of the whole.
-For example, PAR could be replaced with the US Army’s After Action Review (US Army, 1993)—which inspired its basic format in the first place—or by the Feedback Wheel devised by Janet Hurley (Real, 2008).
+workshops (Corneli et al., 2023), and, in the process, PLACARD as a design pattern was elaborated into a design pattern language.
+In this paper, we report on further workshop experiences with PLACARD, and establish a further use case: using the design pattern languages with virtual artificial intelligence (AI) agents.
+Structurally, the paper (§2) provides all necessary background on PLACARD; (§3) outlines our methods; (§4) presents evidence from our recent workshops; (§5) proposes a set of “Design Patterns for Pattern-Competent Agents”; (§6) describes our simple multi-agent workshop simulations; (§7) discusses the study’s limitations, lessons learned, and related work; and (§8) draws conclusions.
+(a) Refer to caption
+(b) Refer to caption
+Figure 1. Text analytics from undetectable.ai
+(a) paper body (b) appendix
+From a methodological standpoint, we drew on OpenAI’s ChatGPT primarily as a research catalyst, and only secondarily as a writing aid (Figure 1(a)), with the aim of exploring the extent to which contemporary chatbot technology can work effectively with design patterns (Figure 1(b)).
+2. Background
+2.1. Motivation
+In an earlier paper, we observed:
+Patterns have previously been discussed in explicitly computational terms—however, that direction of work so far remains mostly at the level of a proposal […], with only limited discipline-specific uptake[.] Could this change, to generalize the kinds of patterns that can be computed with, and make them interoperable?
+(Corneli et al., 2021)
+What was once a speculative trajectory can now be investigated through practical experiments—using generative models to simulate, remix, and refine design pattern languages.
+2.2. Terminology and notation
+We use “pattern” as a shorthand for “design pattern” unless otherwise stated.
+We should acknowledge that our approach to patterns has more in common with the early writing of Alexander (e.g,.
+Alexander and Poyner (1970)) than with the more narrowly technical interpretations found in later software engineering literature.
+Rather than anchoring our work to any single disciplinary tradition, we explore pattern theory from the perspective of design at large, with particular attention to the meta-patterns that shape pattern authoring and use.
+In the process, we aim to address a gap in Alexander’s work.
+He largely presupposes the generative capacity of patterns, without systematically unpacking how they work or under what conditions they foster coherence, value, or transformation.
+Some of our effort will be spent on further exploration of the meaning of “design pattern”; but Leitner’s (2015) definition—“Patterns are shared as complete methodic descriptions intended for practical use by experts and non-experts”—is a good starting point.
+The name “PLACARD” was introduced to denote a remix of three complementary methods for organising collective learning and design (PAR, CLA, and DPL—the acronyms will be expanded and explained just below).
+This anagrammatic device served as a memorable handle in the context of our earlier papers, but it is not at all transparent on first encounter, nor is it quite accurate, since it implies that the pattern can only be implemented using these specific methods, which is not really the case.
+A more transparent and explanatory title for the pattern itself, is Patterns Embody Intelligence.
+The following subsection will briefly retrace the historical development of PLACARD, and introduce “Patterns Embody Intelligence” as a generalized reworking.
+As above, we will use all-caps (like PLACARD) for patterns described in other sources, and small-caps (like Patterns Embody Intelligence) for patterns introduced in this paper.
+2.3. PLACARD — Patterns Embody Intelligence
+The PLACARD pattern was created by aligning three established methods for organising collective intelligence with the typical “context/problem/solution” design pattern schema, as in Table 1.
+The idea here is that the methods can be used to operationalize the creation and use of patterns.
+Table 1. PLACARD components schematized
+Element Method (acronym) What it is
+Context Project Action Review (PAR) A reflective process assessing intentions, actions, and outcomes, useful for learning from experience.
+Problem Causal Layered Analysis (CLA) A futures studies technique analyzing issues across multiple layers—the familiar litany of well-known problems, the systems that underpin them, along with the prevailing worldviews, and deeper myths that constitute the root causes of these problems, along with the possibility of developing alternative perspectives.
+Solution Design Pattern Language (DPL) A framework for capturing and sharing reusable solutions in a structured Context-Problem-Solution format (or similar), further structured by links between such elements.
+This, then, is the PLACARD pattern at its simplest: a somewhat recursive pattern for creating design patterns:
+PLACARD
+Context: Opertating in some context, use the PAR to understand the context.
+Problem: Use the CLA to understand and frame associated problems and their interconnections.
+Solution: Use DPL to elaborate the solution (returning to PAR and CLA as needed).
+That said, stated in this way, PLACARD is perhaps rather less a pattern than a meta-method.
+Our next move was to make the specific methods it mentions into placeholders, and to emphasise instead their functional role.
+PLACARD
+Context: By using PAR (or another sensory method), we are able to identify recurring themes within a context.
+Problem: Then, by using CLA (or another cognitive method), we are able to organize these repeating themes in a structure that exposes the underlying trends, causes, and potential terminating states.
+Solution: With DPL (or another motor method) we can make what we have learned actionable.
+For example, PAR could be replaced with the US Army’s After Action Review (US Army, 1993)—which it itself adapts—or by the Feedback Wheel devised by Janet Hurley (Real, 2008).
 CLA could be replaced by Triple-Loop Learning (Tosey et al., 2012) or Ecological Systems Theory (Bronfenbrenner, 2000).
 DPL could be replaced by Elinor Ostrom’s institutions (Ostrom, 1990), or by a catalogue of literary tropes, and so on.
-Each such selection comes with its own ideological orientation, and typical domain of application.
-2.2. Implementation
-We implemented PLACARD in multiple settings, tailoring our approach to each context:
+A further step is to notice that by the time get to this point PAR, CLA, and DPL are no longer really needed as part of the pattern statement; they are hole-fillers.
+We can properly introduce and start to use new, more general, terminology, better surfacing the underlying pattern (and retitling it) this way:
+Patterns Embody Intelligence
+Context: We operate in a context in which intelligence is available to us. By using a suitable sensory method, we are able to characterize a particular ‘sub-context’ of interest.
+Problem: Then, by using a suitable cognitive method, we are able to organize themes and tensions to identify a ‘problem’ within that sub-context.
+Solution: With a suitable motor method, we can make what we have learned actionable in a way that ‘solves’ the problem.
+In earlier work, we used design patterns to further decompose and operationalize each of these phases.
+We developed patterns for characterizing a context by identifying themes; patterns for organizing structure and exploring tensions; and patterns for transforming insights into action.
+The current paper returns to this familiar terrain with a different aim: to identify and reflect on the patterns that are necessary for working with patterns at all.
+Rather than asking “how might pattern work flourish?”, we now ask “what makes pattern work possible?” The pattern Patterns Embody Intelligence offers a conceptual launch point for this inquiry.
+We will revisit and revise it later in the paper, drawing on insights from our experiments with pattern-capable agents.
+3. Methods
+3.1. Designing and running in-person workshops
+We discussed the basic PLACARD pattern and used it (often in its more adaptable form, PLACARD) to structure workshops in multiple settings, tailoring our approach to each.
+Our hope was to make PLACARD useful to a broader audience, though this was largely unsuccessful, since to our knowledge the authors of this paper are currently the only users of this design pattern.
+Nevertheless, the workshops gave us a core empirical dataset.
 AsianPLoP 2024 (Fujisawa, Japan):
@@
 One challenge discussed was how to be more inclusive of countries across the Asian continent.
-The conference organizers took a positive step by holding the 2025 event in Taiwan, but we cannot state that location decision was influenced by our Focus Group.
-In 2024, onsite attendees helped overcome barriers by clarifying for confused participants in Japanese.
+(The conference organizers took a positive step by holding the 2025 event in Taiwan, but we cannot state that this decision was influenced by our Focus Group.) In 2024, onsite attendees helped overcome barriers by clarifying for confused participants in Japanese.
 This was the only event in this paper conducted in multiple languages.
@@
 A hybrid session titled “AI Future Envisioning with PLACARD” (Tedeschi et al., 2024) with one onsite and three remote facilitators.
-Participants employed the Transfeminist Tech Oracle card game as a creative prompt, then used CLA and PAR in tandem to explore speculative AI futures.
+Participants employed the Transfeminist Tech Oracle card game as a creative prompt, then used CLA and PAR to explore speculative AI futures.
 Due to time constraints and abstraction load, some participants expressed confusion about the method structure, prompting later pattern refinement.
@@
 We learned that we had to explain the rules and definitions prior to playing the game.
+EuroPLoP 2025 (Irsee, Germany):
 A common challenge for all of these workshops is rapid skill acquisition.
-We plan to continue this workshop series, for instance with a EuroPLoP 2025 workshop session.
-As mentioned above, this is a good time to take stock of what we’ve learned so far.
+PLACARD has a lot of moving parts.
+In our EuroPLoP 2025 workshop session, we used the PAR to good effect, but left the other components to one side, replacing didactic instruction with more embodied play.
 3.2. Reflexive Analysis Using the PLACARD Pattern Catalogue
-To analyze the workshops beyond surface-level outcomes, we applied a reflexive methodology grounded in the PLACARD pattern catalogue, presented in PoPII.
-This catalogue is structured into three interlinked categories:
+We drew on AI support to help analyze the workshops.
+We imported the existing PLACARD pattern catalogue into ChatGPT.
+In line with the description above, the catalogue is structured in three categories:
 A. Identifying Themes (“sensory”) B. Organizing Structure (“cognitive”) C. Making It Actionable (“motor”)
-For each workshop, in this paper we use the A/B/C structure to retrospectively assess:
+We retrospectively analysed our write-ups of the above workshops to help assess:
   • Which patterns were implicitly or explicitly activated,
-  • Where breakdowns in understanding or facilitation could be explained by missing or misapplied patterns, and
-  • Where new patterns could be introduced to resolve to unanticipated tensions.
-A brief summary is given in Section 3 of this paper.
-This reflexive approach helped us make sense of divergent workshop dynamics, and the possibility for further patterns that could help make the workshops more rewarding in general.
-For example, Name the Frame was identified in hindsight as something that could have improved onboarding.
-Likewise, Harvest Before You Close was seen as a missing step in several sessions, and this pattern would address missed opportunities for extemporaneous pattern generation.
-Rather than enforcing a rigid rubric, the catalogue functioned as a scaffold for learning from variation.
-It allowed us to compare facilitation dynamics, participant responses, and methodological adaptations across sites—and, crucially, to refine the catalogue itself through lived experience.
-2.4. A Note on Notation
-As above, we will use all-caps (like PLACARD) for patterns introduced in other sources, and small-caps (like Name the Frame) for patterns introduced in this paper, details of which are in the Appendix.
-2.5. Empirical and Design-Based Development of New Patterns
-Across the 2024 and 2025 events, new patterns did not emerge fully formed. Indeed, only one new pattern surfaced and named explicitly during the workshops:
-  • Have Multiple Leaders Onsite – prompted by the hybrid setup at EuroPLoP, where limited physical facilitation led to friction.
-The pattern proposes distributed facilitation roles to support hybrid presence, reduce coordination load, and maintain session energy.
-Other new proto patterns crystallized in our retrospective analysis. For example:
-  • Name the Frame – (a further specification of CONTEXT SETTING from PoPII) this pattern emerged from repeated facilitator observations that participants struggled to orient themselves without a clear method overview.
-  • Borrow the Situation – inferred from participant improvisation in choosing and reinterpreting prompts when predefined situations felt mismatched or abstract.
-  • Start with Sense – (generalizing DÉRIVE COMIX from PoPII) builds on the observation that embodied play or metaphorical framing was often the key to unlocking participation
-In sum, while few patterns emerged as polished artifacts within the workshop frame, the pattern language itself evolved through encounter.
-The workshops provided empirical grounding for evaluating utility, exposing limits, and highlighting latent needs.
-This iterative, design-based process parallels how pattern languages grow in living communities—not through top-down invention, but through grounded, reflective refinement.
+  • Where breakdowns in understanding or facilitation could be explained by missing or misapplied patterns.
+This preliminary analysis was then further developed by hand; the result is presented in Section 4 of this paper.
+We also considered whether new patterns might be introduced that would resolve tensions that had not been anticipated in PoPII; these patterns appear in the Appendix.
 3.3. Developing a comparison case: Working with patterns computationally
 We noticed that it is possible to ask ChatGPT to run simple multi-agent “virtual” workshops.
-These can be compared with the in-person workshops.
-Below are two lightly edited example prompts that were used for setting up these simulations.
-Prompt 1 Given to Generative AI Program I want you to create a multi-agent system with 15 different ChatGPT-based agents.
-I want you to assign each of them the name of a distinct animal, and each of the agents will have some of the typical characteristics of the animal it is named for.
-Then I want you to walk through the 30 “Peeragogical Exercises for a Seminar on Epistemic AI” [elided] in a virtual seminar.
-The final exam is a dialogue between these animals, about the following prompts.
-Please have as much discussion as you’d like among the agents before you give me their final team presentation answering the questions.
-What is epistemic AI?
-How does mathematics relate to this vision?
-What does ‘peeragogy’ mean, and how does it relate to this vision?
-What kind of multi-agent has created this answer?
-Why is this virtual seminar of historical importance?
-Prompt 2 Given to Generative AI Program Remember how we did some simple multi-agent workshops before?
-Let’s simulate the architecture described the paper with some ChatGPT-based agents.
-This time, rather than basing them on animals, we are going to base the agents on a diverse group of Research Fellows in AI.
-Let’s only consider 10 such agents.
-Please include the “Ant” role from the workshop with animals, she will record what happens in which workshop session.
-Please also include the “Dog” role from the animal workshop, who will act as a primary reporter.
-The Fellows will be set the task of designing a series of workshops using the PLACARD method and associated patterns, and the outcome of the series should be a detailed plan of work for a research project on epistemic AI.
-You should do deep research to make sure that the discussions are plausible.
-3. Patterns as Technical, Epistemic, and Reflective Infrastructure
-3.1. Reframing Patterns as Sense-Making Scaffolds
+We ran several such simulated workshops, of which we will discuss two.
+Below is a generic example prompt for creating these light-weight simulations.
+Prompt Template: Multi-Agent Simulation for Collaborative Inquiry Simulate a structured multi-agent workshop involving $N$ distinct agents, each based on a unique [metaphorical persona or professional identity], and endowed with characteristic traits or expertise aligned with that identity.
+The agents should collaborate on a shared task: designing a set of [outputs, e.g., workshops, protocols, strategies] using the [specified framework or method, e.g., PLACARD, Design Patterns, Theory of Change].
+Assign the following roles:
+A Recorder agent who captures the evolving structure and rationale of the discussion.
+A Reporter agent who periodically summarises the current state of collective understanding.
+Optionally, include other roles, such as a Socratic Facilitator agent who prompts clarification, challenges assumptions, and encourages epistemic humility.
+The agents should engage in $M$ discrete sessions, with each session addressing a defined phase or subtask.
+The dialogue should be plausible, reflective of diverse perspectives, and demonstrate progression over time.
+After all sessions are complete, generate a synthetic report or artefact that includes:
+The main findings or decisions of the group
+A reflection on the group process and emergent dynamics
+A meta-level commentary on the kind of multi-agent system that was instantiated
+Please ensure internal consistency and draw on relevant external knowledge or frameworks where applicable.
+4. Patterns as Reflective Infrastructure
 Patterns are often described in instrumental terms: reusable solutions to recurring problems in context.
-As we look into how they work, we see there’s more to it.
-One aspect of this is the use-versus-mention distinction: clearly, a physical alcove is not the same as Christopher Alexander’s “ALCOVES” pattern (Alexander et al., 1977).
-And at both levels, there’s a degree of explicitness.
+However, there’s clearly more to it than that.
+One aspect of this is the use-versus-mention distinction, for example, a physical alcove is not the same as the “ALCOVES” pattern (Alexander et al., 1977).
+There’s also a degree of explicitness.
 When does an indentation become an alcove?
-When does a pattern become part of a language?
-The Start with Sense pattern characterizes various ways of inviting participants to begin inquiry with embodied, perceptual, or intuitive grounding.
-This creates conditions under which problems and possibilities become intelligible.
-Across the A/B/C catalogue, many patterns introduced earlier and from the Appendix operate similarly:
-  • Name the Frame helps surface the underlying structure of an activity or method.
-  • Borrow the Situation allows participants to reinterpret or co-create the context of inquiry.
-  • Meta-Reflection Loop supports collective understanding of what just happened and why.
-These are cognitive footholds, ways of managing ambiguity and distributing attention. They help groups compose a shared reality, especially when dealing with complex or unfamiliar material.
-From this perspective, the value of a pattern lies not just in its repeatability but in its sensitivity to situated meaning-making.
-A good pattern doesn’t just say “do this”; it says “look here, attend to this tension, notice this affordance.” It invites participants to see the world in terms of actionable differences.
-Reframing patterns this way brings PLACARD into closer dialogue with traditions such as:
-  • Phenomenology—emphasis on situated perception (Moran, 2000)
-  • Participatory design—emphasis on co-articulation of problems (Spinuzzi, 2005)
-  • Grounded theory—emphasis on emergent categorization (Glaser, 1992)
-It also prepares the ground for computational application.
-For artificial agents, much as for humans, sense-making begins with activation of relevance—the ability to determine what matters in a given context.
-In PLACARD, patterns serve as activation scaffolds: semi-formal cues that guide what should be noticed, named, or negotiated.
-Each of its layers—A/B/C—corresponds to a distinct mode of learning.
-Crucially, design patterns—or, indeed, any substitute schema that plays the same role—achieve the function of linking these modes, enabling transitions between experience, inquiry, and abstraction.
-We propose that patterns support at least three interdependent forms of learning:
-  1. (1)
-    Reflective Learning (Boyd and Fales, 1983): Patterns such as PROJECT ACTION REVIEW  and Meta-Reflection Loop support meaning-making after action.
-They help participants surface experiences, identify what shifted, and consider next steps.
-This mode is retrospective and grounded in collective memory and sense-making.
-  2. (2)
-    Design Learning (Burdick, 2012): Patterns like Borrow the Situation or Pattern the Play create scaffolds for reimagining problems, proposing alternatives, and testing ideas.
-Design learning is generative, situated in ambiguity.
-Patterns here serve as framing devices that guide creativity while constraining chaos.
-  3. (3)
-    Model Learning: Patterns like Name the Frame and Harvest Before You Close deploy abstraction—capturing structural insights from activity.
-This form of learning supports knowledge representation and the translation of local insights into portable conceptual tools (e.g., (Ariyaratne, 1977)).
-Patterns enable movement between these modes. For example:
-  • A reflection may generate a design insight that prompts reframing.
-  • A design intervention may require formal modeling to share or extend.
-  • A model may provoke a return to embodied or emotional grounding.
-3.2. Summary of the evidence base from our in-person workshops
-Table LABEL:workshop-summary summarizes the use of PoPII patterns in our workshops. Further details are the Appendix.
-Table 1. Summary of patterns used in workshops and potential for improvement
-Category/Patterns Used Well Needs Refinement
-A./DÉRIVE COMIX, DO YOUR RESEARCH, CONTEXT SETTING, PAR Overrun — PAR provides shared cohesion, when used; our PAR-inspired survey has a different feel.
-Context-setting could also be more cohesive: Name the Frame.
-Explore more creative, embodied PAR.
-B./GOING META, FACILITATOR and FUNCTIONAL ROLES Mixed — time + onboarding were limits.
-Introduce meta-language tools earlier and more clearly.
-Test whether embodied cues help people get to worldview/myth.
-C./REINFUSE EXPERTISE Implicit — patterns used, not made. Stronger follow-through on actionable steps; Harvest Before You Close. Explicitly support pattern emergence from cards/music activities.
+When does a proto-pattern become a pattern?
+One typical answer to the latter question is that the pattern should be applicable in suitably varied contexts.
+A further implication is that a pattern might be “observed” empirically even when it has not been formally “applied” in a design sense.
+After all, alcoves existed long before 1977!
+Although we did not explicitly structure our recent series of workshops using PLACARD pattern language (rather, we used the earlier PLACARD pattern), nevertheless the workshops continue the spirit of that work, and the patterns themselves could often be rationally reconstructed from the evidence base that the current round of workshops provides.
+Details of the previous pattern statements are to be found in (Corneli et al., 2023).
+The qualitative analysis here compiles evidence of their implicit use, or the use of a similar strategy, looking across the in-person workshops mentioned above.
+This is not intended to be a ‘feature grid’ showing which workshops used which patterns.
+A. Identifying Themes
+DÉRIVE COMIX
+Application: In several of the workshops, participants used Transfeminist Tech Oracle cards (with ‘suits’: object, value, body/territory, situation) as interpretive prompts to spark imagination—unfamiliar juxtapositions mirroring the urban dérive or imagined stroll in a more conceptual sense.
+Observation: The card game facilitated divergent thinking akin to exploring unfamiliar terrain.
+In the 2025 EuroPLoP workshop, we did a similar exploration, using less complex images from the card game Dixit.
+CONTEXT SETTING
+Application: Early slides and explanations introduced key concepts (e.g., design patterns, CLA).
+Observation: Participant feedback highlighted a need for clearer up-front orientation.
+Improved alignment with this pattern could increase accessibility.
+That said, in the 2025 EuroPLoP workshop, observed that some “interface friction” can sometimes be a benefit for workshop outcomes, even if it is frustrating for participants.
+SHARE BACK
+Application: When there were a sufficient number of attendees, they did break into small groups and share back.
+DO YOUR RESEARCH
+Application: Past workshop experiences were integrated into the design; earlier in-house uses of CLA and PAR informed subsequent sessions.
+Example: Reference to Anticipation 2019 and prior events illustrates “pre-research” informing present design.
+B. Organizing Structure
+GOING META
+Application: The overall project itself exemplifies recursive methodology—using PLACARD to better understand and refine PLACARD.
+Example: In two of the workshops, participants were prompted to reflect on and imagine both the methodology and the subject (AI futures).
+PATTERN LANGUAGE COMPONENTS
+Weak application: The workshops in this series typically focused on the PLACARD components (CLA + DPL + PAR).
+We typically did not go into further detail of DPLs.
+While it is true that CLA, DPL, and PAR do suggest a way to think about pattern language components, because they are complex, they are not always user friendly.
+Recommendation: In some of the workshops reported on in our previous paper, we used light-weight keywords (“however”, “because”, “specifically”) to scaffold understanding of design pattern internals incrementally.
+In the current series of workshops, that might have been less useful in any case, because attendees were already familiar with pattern languages.
+We had supposed that, within this population, the Transfeminist Tech Oracle cards would prompt pattern-like thinking, but we did not assess whether it did or not.
+FACILITATOR ROLES
+Application: Hybrid facilitation with distinct roles (onsite leader, remote supporters) distributed facilitation and modelled different levels of participation. (Cf. FUNCTIONAL ROLES below.)
+C. Making it Actionable
+THE FUTURE BEGINS NOW
+Weak application: Using the card game, participants generated imagined futures.
+Recommendation: While creative ideas about the future emerged, follow-through actions or small-scale pilots were typically not described.
+In the 2025 EuroPLoP workshop, and other contexts where a PAR was carried out, “next steps” were explicitly surfaced.
+These were often next steps for the workshop leaders, not workshop participants, but in the 2025 EuroPLoP session, we explicitly directed participants to think about potential next steps for themselves.
+FUNCTIONAL ROLES
+Weak application: We did not explicitly involve workshop participants in any formal roles.
+(In the 2025 EuroPLoP workshop, participants did have the opportunity to take on an untitled light-weight roles, such as introducing a simple musical theme in a circle music exercise.) Recommendation: Our earlier work developed several explicit roles for participants to explore: these or other roles could be explained and offered, even if they are not taken up—however, we must acknowledge that there is a time-cost and cognitive load when doing so.
+PROJECT ACTION REVIEW (PAR)
+Application: Several sessions used the PAR pattern at the end to reflect, share learnings, and gather insights.
+In some cases where the PAR was not used, we shared a survey with questions informed by the PAR.
+Recommendation: An in-person PAR and a survey are very different experiences for participants.
+REINFUSE EXPERTISE
+Application: Participants brought diverse professional and cultural perspectives.
+Their expertise was implicitly integrated into the discussion in shared reflection on topics such as AI futures and governance.
+Recommendation: A more explicit arc—temporarily setting aside expertise to foster common ground—is more in the spirit of this pattern.
+In the 2025 EuroPLoP session, participant expertise was explicitly invited in the concluding PAR.
 5. Design Patterns for Pattern-Competent Agents
-The PLACARD methodology, as developed and refined across multiple participatory workshops, demonstrates the power of design patterns not only as tools for structuring action, but also as scaffolds for shared sense-making and reflection.
-Building on this foundation, we now consider a forward-looking design space: the development of pattern-competent artificial agents capable of engaging with design patterns as part of collaborative inquiry.
+The typical use of design patterns is as methods for structuring action.
+As the previous section illustrates, they can also be used to scaffold sense-making and reflection after the fact.
+Building on these foundations, we now consider a forward-looking design space: the development of pattern-competent artificial agents capable of engaging with design patterns in a process of collaborative inquiry.
 These agents would be expected to:
   • Recognize situations where a pattern might apply,
-  • Use flexiformal representations of pattern logic,
-  • Operate as a collective, i.e., to form multi-agent systems (MAS)
+  • Use suitably flexible representations of pattern logic,
   • Participate in reflection loops that evolve the infrastructure itself.
-4.1. Recognizing Pattern-Relevant Situations
+  • Operate as a collective, i.e., agents combine in multi-agent systems (MAS),
+Our existing patterns give a start in this direction, at least insofar as they could help us to imagine collaborative workshops that include or are populated by artificial agents.
+However, the aim of embedding the abilities mentioned above within the computational agents will suggest some new patterns.
+First, let’s consider the above points in turn.
+5.1. Overview of requirements
+5.1.1. Recognizing and Responding to Pattern-Relevant Situations
 The capacity to recognize situations where a pattern might apply lies at the heart of pattern competence.
@@
 It requires the ability to perceive and interpret contextual cues, to identify when a recurring structure or problem-type is emerging, and to discern salience: which aspects of a situation matter, and why.
-In participatory settings, such recognition is informed by experience, tacit knowledge, and social interaction.
+In typical participatory settings, such recognition is informed by experience, tacit knowledge, and social interaction.
+For people who are used to working with design patterns, a new experience might be thought with some light attention going to an existing pattern catalogue, without being constrained to only think in terms of known patterns.
 When developing artificial agents, the challenge is to develop mechanisms—symbolic, statistical, or hybrid—that allow for similarly situated recognition.
-Here’s an example showing how this might work for today’s computer systems:
-Example: A chatbot trained to handle customer service inquiries recognizes that a user is “frustrated,” even if the words don’t match prior examples exactly—but the tone, pacing, and word choice are similar to previously tagged “frustrated” interactions.
-Mechanism: A statistical similarity model (e.g., embedding distance + fine-tuned classifier) matches this new situation to prior ones and adjusts the agent’s response accordingly (e.g., a more empathetic tone or an escalation offer).
-Next-generation systems might deal with the same scenario in more flexible ways, for instance, not only on the basis of stimulus-response pairs, but more fluidly—for instance, understanding how it was that the customer became frustrated and initiating changes to the company policy accordingly.
-This more meaningful form of ‘pattern recognition’ requires readiness to see and make sense of structure, and to reframe experience in meaningful, actionable terms — rather than to set data aside as a block, or (which is much the same) to cope with it in a purely reactive way.
-4.2. Flexiformal Representation of Pattern Logic
-To support agent-based reasoning, patterns must be expressed in a format that is both structured and open-ended.
-Drawing on the modified pattern structure used in PLACARD—ContextIf … (But)… ThenBecause…—each clause can be understood as a layer of scoping and activation:
+Here’s an example showing how this sort of process works for today’s computer systems:
+Example: A chatbot trained to handle customer service inquiries recognizes that a user is “frustrated,” even if the words don’t match prior examples exactly—but the tone, pacing, and word choice are similar to previously interactions that were tagged “frustrated”.
+Mechanism: A statistical similarity model (e.g., embedding distance + fine-tuned classifier) matches this new situation to prior ones and adjusts the agent’s response accordingly (e.g., to adopt a more empathetic tone or make an escalation offer).
+Next-generation systems might deal with the same scenario in more flexible ways, for instance, by attempting to build a narrative understanding of how it was that the customer became frustrated, and proposing changes to the company policy that would avoid such problems in the future.
+This more meaningful form of ‘pattern recognition’—which is set up in a way that supports ‘pattern application’—requires readiness to see and make sense of structure, and to reframe experience in appropriately meaningful, actionable terms.
+5.1.2. Flexiformal Representation of Pattern Logic
+The primary difference between design patterns and computer programs is that design patterns are both structured and significantly open-ended.
+Their structure comes primarily from the use of a consistent schema (e.g., context, problem, solution), and from links between patterns.
+Their open-endedness lies in their reliance on user interpretation and situational judgment.
+Insofar as they ‘operate’ at all, patterns do so through enactment in specific contexts, rather than via deterministic execution.
+Nevertheless, as we explore further here—and as been discussed by various authors before (Corneli et al.
+(2018) and references therein, particularly to the work of T.
+P.
+Moran and Rivka Oxman)—patterns can also be framed in computational terms.
+Drawing on the adapted pattern template we used to describe our PLACARD patterns—ContextIf … (But)… ThenBecause…—each clause can be understood as a layer of scoping and activation:
 Context::
-a broad frame (task domain, discourse setting) that activates relevant subsystems,
+A broad frame—task domain, discourse setting, data ingest—that activates relevant subsystems.
 If condition::
-a salient event or state within that frame,
+A salient event or state within that frame, which can be selected by preliminary processing, such as filtering or querying.
 But::
-a tension or constraint that introduces conflict,
+A tension, constraint or conflict—established through additional filtering, querying, data alignment—which suggests the possibility of, or need for, further processing.
 Then::
@@
 Because::
-While logic and reasoning have been core topics in computer science from its outset (Sowa, 1999), further research is needed to determine what sorts of reasoning—clearly combining abductive, deductive, inductive, and analogical modes—and what types of logic—classical, fuzzy, modal, monotonic, non-monotonic—are used by humans to think about patterns.
+A set of critical checks or tests to take before selecting the action; typically framed with respect to some existing body of empirical evidence; possibly also useful for supplying auditable reasoning alongside the action.
 For now, we refer to the overall approach as “flexiformal” (Kohlhase, 2012), because this leaves the door open to combining different formal systems in a suitably flexible fashion.
-Simply adding more patterns won’t be enough to implement human-like reasoning in pattern-capable systems.
-For one thing, the patterns need to be used to dynamically allocate cognitive resources.
-4.3. Reflection Loops and the Evolution of Epistemic Infrastructure
-To sustain meaningful pattern use, agents must also be able to reflect.
-This means embedding meta-level structures like the PAR into the environment itself.
-Reflection loops should support the capture of learning, the revision of pattern catalogues, and the emergence of new design insights.
-In future architectures, agents and humans alike can annotate, revise, and remix patterns, and participate in the shared cultivation of ‘living’ knowledge systems, in which patterns are part of a generative infrastructure for collective intelligence.
-We’ve previously imagined patterns being used within a “Discovery Machine” (Tedeschi et al., 2024).
+Logic and reasoning have been core topics in computer science from its outset (Sowa, 1999); however further research would needed to determine what sorts of reasoning—pesumably, combining abductive, deductive, inductive, and analogical modes—and what types of logic—classical, fuzzy, modal, monotonic, non-monotonic—are used by humans when they think about patterns.
+Furthermore, simply specifying more patterns won’t be enough to implement human-like reasoning in pattern-capable systems.
+As we’ve indicated here, a key thing that the patterns need to do is to dynamically allocating cognitive resources.
+The patterns’ internal structure and external relationships with each other (and with other interfaces) constrain the system’s ability in that regard.
+5.1.3. Reflection Loops and the Evolution of Epistemic Infrastructure
+To sustain meaningful pattern use, agents would have to be able to reflect; noting new patterns, revising them, including changing the way they relate, and deprecating or deemphasising patterns that are no longer relevant.
+This requires embedding meta-level review structures—perhaps something like the PAR—within the agents or their interactions.
+In future architectures, agents and humans alike might work together to annotate, revise, and remix patterns, and participate in the shared cultivation of ‘living’ knowledge systems, in which patterns are part of a generative infrastructure for collective intelligence.
+(We’ve previously imagined patterns being used within a “Discovery Machine” (Tedeschi et al., 2024).)
+5.1.4. Environments in which pattern-capable agents can form pattern-competent MAS
+Relevant multi-agent system (MAS) environmental features would include scaffolding for the social and cognitive dynamics of inquiry, such as coordination, negotiation, and reflection (as touched on above).
+Such environments have the capacity to support multiple, interacting Baldwin effects—individual learning loops that influence each other over time as the agents and their interactions evolve (Red’ko et al., 2005).
+Rather than simply evaluating success in terms of task completion or efficiency, pattern-competent MAS environments might invite questions about “how” solutions are created.
 5.2. Rethinking the Pattern Catalogue
-The PLACARD A/B/C pattern catalogue was created with human participants in mind.
-It emphasizes social affordances, embodied prompts, and facilitation moves.
-While some of these translate readily into multi-agent contexts—e.g., “Meta-Reflection Loop” or “Pattern the Play”—others rely on human sensory input or informal framing strategies that are not directly available to AI systems.
-Thus, we propose a tripartite model of pattern types needed for pattern-competent agents, generalizing the layers we used previously.
-1. Interactional Patterns (generalizing the sensory layer, e.g., PAR)
-PAR structures collaborative reflection and coordination.
-Interactional patterns in general guide agents in playing roles, reflecting, and negotiating meaning in collaborative settings—essentially the micro-level behaviors that mirror PAR’s inquiry loop.
-Examples include:
-  • Role Reveal: Adopt explicit roles in multi-agent dialogue.
+The series of short reflections above suggests an extension to the way we were thinking about the PLACARD pattern.
+PLACARD was cast along classic psychological lines (Hilgard, 1980)—and the sensory/cognitive/motor facilities have also been taken over by roboticists (Cassimatis et al., 2004).
+However, it does not specifically talk about the environment in which the agents it considers would be situated.
+We would expect an undersea robot and a Mars rover to have very different functionality, even if they have the same theoretical underpinning; and, perhaps even more to the point, we would expect a robot that operates as part of a swarm to work very differently from one that operates individually.
+Perhaps because the default assumption with PLACARD was that the actors were humans operating in standard human ways, it did not deal with these contextual issues in any detail.
+Accordingly, the four dimensions outlined above suggests that a pattern language for pattern capable agents should be extended—retaining the A.
+sensory, B.
+cognitive, and C.
+motor segments, and adding a fourth segment, Denvironmental or developmental patterns.
+With this reflection in mind, we propose a quadripartite model of pattern types needed for pattern-competent agents, generalizing and extending the layers we used previously, and giving further expression to the requirements outlined above.
+Some examples follow (detailed pattern statements are in the Appendix).
+A. Interactional Patterns (generalizing the sensory layer, e.g., PAR)
+Interactional patterns in general guide agents in noticing when patterns can be applied, playing roles, reflecting, and negotiating meaning in collaborative settings.
+To a reasonable extent these behaviors mirror PAR’s inquiry loop, but instead of being a stand-alone reflection ritual, they would describe a broader set of practices of mutuality within an environment.
+Examples could include.
   • Meta-Reflection Loop: Periodically engage in structured self-review.
   • Timebox the Core: Manage resource allocation and attention windows.
-These patterns require some translation to make it understandable to the machine, e.g., encoding roles as functions in MAS architectures or structuring turn-taking rules.
-2. Cognitive Patterns (expanding upon the existing cognitive layer, e.g., CLA)
-CLA unpacks multiple layers of meaning—litany, system, worldview, myth.
+  • Role Reveal: Adopt and express explicit roles to guide multi-agent dialogue.
+B. Cognitive Patterns (expanding upon the existing cognitive layer, e.g., CLA)
 These cognitive patterns help agents reason across contexts and detect tensions, embedding depth and nuance in internal reasoning.
-They are interpretive structures, supporting internal reasoning and pattern use:
+Whereas CLA unpacks multiple layers of meaning, these patterns describe further interpretive structures, supporting internal reasoning and pattern use:
   • Pattern Activation: Given a context and goal, retrieve applicable patterns.
@@
   • Candidate Move Generation: Translate pattern suggestions into operational plans.
-3. Infrastructural Patterns (generalizing the motor layer, e.g., DPL)
-DPL captures reusable structures for action.
-These infrastructural patterns define persistent templates and system-wide protocols that support reuse, feedback, and evolution of the broader system design.
-For example these are 3 meta-patterns that could enable engagement with patterns:
-  • Shared Memory Architecture: Persistent storage of pattern use and modification.
-Putting patterns where people can get them, like the catalogue, akin to the SCRAPBOOK from Patterns of Peeragogy (Corneli et al., 2015)
+C. Infrastructural Patterns (generalizing the motor layer, e.g., DPL)
+These infrastructural patterns define templates and system-wide protocols that support various forms of interaction with and through the broader system—reuse, feedback, and evolution of the broader system design.
+Whereas DPL captures general structures for action, these patterns focus on the self-coherence implied by ‘action’, insofar as it always involves an actor.
+For example, we begin here with three meta-patterns that would enable agents to engage with a collection patterns:
+  • Shared Memory Architecture: Persistent storage of pattern use and modification, putting patterns that aren’t being actively used someplace where they can be retrieved (akin to the SCRAPBOOK from Patterns of Peeragogy (Corneli et al., 2015))
   • Feedback Rhythms: Protocols for scheduling reflection and pattern revision.
   • Pattern Diffusion: Mechanisms for propagating new or adapted patterns across agent collectives.
-They govern the architecture of the agent system much like DPL governs the architecture of human practice.
-4.5. Desiging Environments for pattern-competent MAS
-In addition to broader contextual information, relevant multi-agent system (MAS) environments would include the social and cognitive dynamics of inquiry: coordination, negotiation, reflection.
-Such environments have the capacity to support multiple, interacting Baldwin effects—individual learning loops that influence each other over time as the agents and their interactions evolve (Red’ko et al., 2005).
-This distributed adaptation mirrors cultural evolution: strategies become shared, revised, and recontextualized.
-Rather than simply evaluating success in terms of task completion or efficiency, MAS environments invite questions about “how” solutions are created.
-This reflection constitutes suggests a key amendment to PLACARD.
-Whereas PLACARD was described in classic psychological terms (Hilgard, 1980) — which are also taken over by roboticists (Cassimatis et al., 2004) — it does not specifically talk about the environment in which the agents it considers would be situated.
-In robot terms, for example, we would expect an undersea robot and a Mars rover to have very different functionality, even if they have the same theoretical underpinning.
-This way of thinking suggests that the PLACARD pattern language should be extended to be an “A/B/C/D catalogue”, where the fourth section contains environmental or developmental patterns.
-With this in mind, in addition to further detail on the candidate patterns for pattern-competent agents — as outlined above — the appendix also includes some patterns that would guide the design of their working environments.
-5. Virtual Multi-Agent Workshops in the Spirit of PLACARD
-To explore the idea of pattern-competent agents developed Section 4, we conducted a series of virtual workshops using fictional agents within ChatGPT.
-Using the PLACARD method, those agents rotated through breakout sessions, engaged in pattern-guided inquiry, and made use of a curated and emergent pattern catalogue.
-It is worth emphasizing that the workshops were computer-aided design fictions (we have previously explored this way of working “by hand” (Danoff et al., 2019), and we’re aware that fictional works can have unexpected real-world ramifications, e.g., (Wikipedia contributors, 2025)).
-ChatGPT can simulate plausible dialogues, behaviors, and reflective loops based on design patterns, but it cannot build and run custom agent architectures.
-All ‘agents’ were textual personas that were simulated by one model.
-The agents were plausibly able to discuss the design patterns and employ them in light-weight design exercises.
-Below is an excerpt from one such workshop, in which several virtual agents begin collaboratively applying the PLACARD methodology to plan a research project on epistemic AI.
-The agents have somewhat stereotyped fictional biographies which to some extent shape their responses (e.g., “Carlos – Epistemology & Systems Philosophy”; “Mei – Participatory Design & Feminist STS”).
-Excerpt: PLACARD Planning Dialogue on Epistemic AI Carlos: I propose we begin with CONTEXT SETTING.
-If we’re designing a research project on epistemic AI, we need a shared sense of what “epistemic” means in our context.
-Priya: Agreed.
-For me, it’s about how agents—human or artificial—come to know, not just calculate.
-So I’d emphasize design patterns that support interpretive, situated reasoning.
-Raj: Right.
-Maybe we could use Borrow the Situation—look at real-world cases of failure or success in knowledge systems.
-Misinformation management?
-Scientific discovery pipelines?
-Sofia: I think we need to Name the Frame explicitly.
-Are we researching tools for knowledge validation?
-Or institutional models for human-AI epistemic partnerships?
-Mei: Maybe both.
-What if we used Pattern the Play to design speculative workshop sessions?
-Each session could simulate a different epistemic tension—uncertainty, disagreement, novelty—and trial agent responses.
-Yuki: And we need some kind of Meta-Reflection Loop.
-A structure where, after each workshop, we reflect not just on what we learned, but how the patterns performed.
-This could feed into our developmental arc.
-Carlos: So the goal of the series isn’t just content generation—it’s to test, refine, and expand the pattern catalogue itself?
-Sofia: Exactly.
-And that aligns with the DO YOUR RESEARCH pattern: we’re generating grounded material and modeling the tools that structure knowledge formation.
-Raj: We’ll need a pattern-aware toolchain too—maybe a Pattern Knowledge Repository that evolves with each workshop.
-Otherwise we lose coherence.
-Priya: Let’s add a Timebox the Core phase to every session.
-This gives us just enough structure to avoid sprawling, without over-engineering.
-Mei: I’ll draft a plan using PLACARD as a scaffolding template.
-We can each take ownership of one layer—PAR, CLA, DPL—and develop sub-patterns and protocols that fit.
+D. Environmental or Developmental patterns (e.g., institutional arrangements)
+These patterns help multi-agent systems evolve over time and maintain coherence across shifting contexts.
+They provide orientation at the level of institutions, ecological awareness, and collective change.
+  • Institutional Drift: When pattern use becomes routine but out of sync with current challenges, trigger a meta-level review to restore relevance.
+  • Norm Seed: At the start of a new MAS, plant simple, precedent-based mini-patterns to guide initial interactions and shape future dynamics.
+  • Boundary Oscillation: When divergent roles or norms emerge across domains, activate a boundary negotiation process to realign shared pattern use.
+  • Reflection Scaffold Upgrade: If reflection fails to generate insight, introduce new layered tools or protocols to support deeper system learning.
+6. Virtual Multi-Agent Workshops with Patterns
+To explore the idea of pattern-competent agents developed in Section 5, we conducted a series of virtual workshops using fictional agents within ChatGPT; we report on two such workshops here.
+6.1. Simulated PLACARD workshop
+Agents were invited to use the PAR, CLA, and DPL methods, and were given access to existing patterns from our previous work with PLACARD, along with a further collection of new patterns covering similar topics (Appendix:(1)).
+The prompt for the workshop appears in Figure 2, edited for readability.
+This adapts the prompt template from Section 3.3.
+The agents that were created were given somewhat stereotyped fictional biographies, which to seem to shape their contributions to the dialogue.
+However, ChatGPT cannot build and run custom agent architectures.
+All ‘agents’ were textual personae that were simulated by one model.
+Virtual workshops in this style should be understood from the outset as computer-aided design fictions.
+ChatGPT can simulate plausible dialogues, behaviors, and reflective loops based on design patterns.
+In this example, it would be most accurate to say that a virtual workshop was “simulated”, rather that that it “actually happened”.
+Some key experts from the ChatGPT’s trace (obtained via the Think Deeply feature) are as follows:
+I’m gathering key workshop patterns and considering roles like “Functional Roles” and “Facilitator Roles” to ensure synergy and leverage interdisciplinary insights.
+Wondering about talents like Omar, focusing on knowledge representation, and Yuki, adding a creative vision to multi-agent interaction. Considering resources and striking a balance.
+OK, let’s see. Participants are thinking through research questions in breakout rooms, covering math, multi-agent, and meta topics, ensuring a balanced mix of expertise.
+Participants are circling back old and new patterns, making iterations, and planning the final steps. Pattern names stand out, guiding their approach for a comprehensive strategy.
+See Figure 3 for a summary of the workflow, and main research questions that were considered.
+This figure summarises a considerably longer (22-page) workshop report that was generated at the end of the run.
+[^1] https://bit.ly/virtual-placard-workshop.
+The linked supplemental information includes the verbatim prompt which is glossed in Figure 1.
+Relative to our discussion of patterns: looking over the report, existing patterns (like REINFUSE EXPERTISE, GOING META, Start with Sense, Name the Frame, etc.) have been applied in a plausable way.
+Some new proto-patterns e.g., (“Lab Shadowing” and “Checklist Manifesto for AI”) were also proposed.
+Note that the high-level themes in the workflow—the breakout rooms—are derived from earlier chats with ChatGPT.
+While the methods we have described are repeatable, but it would probably be impossible to reproduce the experiment exactly, even using the same ChatGPT account (particularly because other chats have taken place using that account in the mean time).
+ChatGPT is able to generate plausable dialogue snippets on request.
+If it had asked for “a detailed simulation of the breakout session” as an output, or immediately after running the workshop simulation, it would likely produce something plausible.
+In the following section, we did just that.
+6.2. Simulated pattern-capable agents workshop
+Reflecting on the strengths and shortcomings of the virtual workshop described above, we tried another design.
+The prompt is given in Figure 4.
+This time, the collection of patterns for pattern-capable agents and environments (in Appendix(2), Appendix(3)) played a key role in the prompt.
+6.2.1. First round
+With Figure 4 as the overall framing, the subprompt and short answer for this round was:
+Q1. “How would we characterize, in prose, the constraints and features that the agent-environment patterns supply for collective problem solving?”
+A1.
+“Each pattern-agent imposes certain constraints that structure the collective problem-solving process (preventing particular failures or excesses), and each provides features that enable or enhance the group’s ability to solve problems together.”
+Full outputs are available as a supplement. [^2] https://bit.ly/virtual-agents-workshop-i The following changes to the pattern catalogue were proposed on the way to the short answer above.
+EVOLVE – Timebox the Core: The Reflect in Layers and Timebox the Core agents collaboratively proposed a change to the Timebox the Core pattern.
+The pattern was evolved to include a provision for deeper reflection: when core tasks are finished or when a deeper insight is flagged during the session, Timebox will allow scheduling of a follow-up for that deeper exploration rather than strictly ending the session.
+This evolution was agreed upon by the group and validated via Legitimate Iteration for consensus and proper documentation.
+GENERATE – Outcome Synthesizer: The agents identified a gap – no existing pattern solely focused on integrating multiple contributions into a coherent solution.
+Meta-Reflection Loop (supported by Candidate Move Generation and others) proposed a new pattern-agent, tentatively named Outcome Synthesizer.
+With encouragement from Self-Patterning Mandate and formal group approval (again facilitated by Legitimate Iteration), this new pattern was introduced.
+Outcome Synthesizer is designed to gather and synthesize the group’s disparate inputs into a unified outcome or summary, ensuring the collective output is coherent.
+(Note that the Outcome Synthesizer is effectively a rediscovery of the WRAPPER peeragogy pattern (Corneli et al., 2015)).)
+6.2.2. Second round
+This time, we used the following subprompt, and received the accompanying short answer: [^3] https://bit.ly/virtual-agents-workshop-ii
+Q2.
+“How would the pattern-capable agents take over the work from the previous PLACARD workshop on epistemic AI, as captured in a printout describing the same, and build on it?
+Please phrase the answer as a prompt for another round of agent-based simulation.”
+A2. “How might we run a workshop where AI Fellows co-develop new institutional forms for epistemic AI research, supported by our pattern-agent infrastructure?”
+In addition to “breaking the fourth wall” and proposing to run a workshop with other humans (or simulated humans?), something else interesting and a bit strange happened in the dialogue itself.
+The agent workshop started to treat some of the PLACARD patterns that the system had encountered previously as ‘agents’, effectively drawing them into the dialogue.
+One outcome of the workshop was to propose an evolution to Increase Participant Control—even though this was not one of the patterns that had been submitted for revision.
+While the answer A2 could be interesting to adapt and explore in a further workshop—and it is indeed quite intriguing to think about setting up a hybrid workshop with some agents and some humans via a chat interface [^4] https://github.com/IngoKl/PolyBotConversation—it is perhaps best to close the curtain on this series of AI workshops, and discuss what we have seen so far.
+“Remember how we did some simple multi-agent workshops before?
+Let’s simulate the architecture described in our paper-in-progress with some ChatGPT-based agents.
+We are going to base the agents on a diverse group of Research Fellows in AI.
+Let’s consider 10 such agents.
+Please additionally include a Recorder who will keep track of what happens in which workshop session.
+Please also include a Reporter who can speak to what happens.
+The Fellows will be set the task of designing a series of workshops using the PLACARD method and associated patterns, which are available in the file patterns.txt.
+The outcome of the series should be a detailed plan of work for a research project on epistemic AI.
+You should do deep research to make sure that the discussions are plausible.”
+Figure 2. Prompt Given to Generative AI Program to create and run a Virtual Workshop (lightly edited for readability)
+Round 1: Exploring Key Research Questions (Ideation Round)
+  Breakout room 1 (Epistemic Modeling & Validation in Mathematical Domains): (Q1) What frameworks can rigorously verify AI-generated proofs and integrate with human mathematicians’ validation workflows?
+and (Q2) How can we model the epistemic confidence in a mathematical claim across human–AI collaborations?
+  Breakout room 2 (Collaborative Design Patterns for Multi-Agent Systems): (Q3) What design patterns can facilitate transparent communication and division of labor in mixed teams of humans and AI agents?
+and (Q4) Which collaborative strategies best mitigate conflicts or errors in autonomous multi-agent research systems?
+  Breakout room 3 (Metascientific Observation of AI Research Practice): (Q5) What methods can we use to observe and measure current AI research practices (in labs and publications) to identify gaps in transparency and collaboration?
+and (Q6) Which interventions (patterns, policies, or tools) could improve reproducibility and inclusivity in AI research communities?
+Coffee break
+  “Early cross-pollination: Fellows are already trading metaphors across domains – a sign of interdisciplinary synergy.” PAR
+  “I realized my ‘big question’ about AI proofs was actually two questions once I considered both technical and human angles.”
+Round 2: Developing Approaches and Patterns (Design Round)
+PAR “We’re actually developing our own pattern for how to run a project like this – maybe we should write that down too.” SHARE BACK plenary
+Round 3: Integration and Planning (Convergence Round)
+Integration and Closing Discussion
+Figure 3. Reported workshop structure and selected quotes
+You have been given 20 design patterns describing pattern-capable agents and pattern-capable agent environments.
+I want you to run a multi-agent textual simulation where each pattern is an agent.
+“Agents” will operate in an environment which has the agent-environment patterns as constraints or features.
+You are tasked with the following problem:
+We need to design a multi-agent protocol that can address real-world problems.
+There are 20 of us now in total.
+We can create more agents as needed.
+Each ‘agent’ is a pattern, and each pattern is an ‘agent’ — including the agent-environment patterns.
+Together, we are to generate sub-problems that elaborate this challenge and work through them in a self-organised way.
+In addition to “DISCUSS”, our two main moves are:
+  • EVOLVE: Revise or adapt the description of an existing pattern (including changing its trigger conditions, actions, or connections).
+  • GENERATE: Create a new pattern-agent who will join the population, based on observed needs or tensions.
+You may also EVOLVE or GENERATE agent-environment patterns.
+These define the operational constraints or affordances within which all agents must function.
+Changes to these patterns will be treated as systemic shifts, impacting the context for subsequent rounds.
+We will, interactively, run several rounds of simulation. Each round will consist of:
+  1. (1)
+    Agreeing upon a sub-prompt or tension for the group to address.
+  2. (2)
+    Discussion between the agents.
+  3. (3)
+    One or more agents making EVOLVE or GENERATE moves in response.
+  4. (4)
+    Optional meta-commentary or reflection from other agents.
+At the end, output a heredity tree showing how the agent population evolved — i.e., which agents were changed, what was added, and what led to each transformation.
+Please include:
+  • A summary of the dialogue for each round
+  • A clear record of EVOLVE / GENERATE moves
+  • The updated pattern descriptions (for any evolved or new agents)
+  • A final heredity tree visualising the evolution
+Figure 4. Prompt for evolving the patterns for pattern-capable agents and agent-environments
 7. Discussion
 7.1. Limitations and Lessons Learned
-While the outcomes of our PLACARD-based workshops (both in-person and simulated) were promising, several limitations emerged that highlight areas for refinement—both in the method itself and in our implementation of it.
-Generally, these limitations correspond to lessons that could be applied in future work.
+Although the outcomes of our workshops — both in-person and simulated — were promising, several limitations emerged that highlight areas for refinement, both in the method itself and in our implementation of it.
+Generally, these limitations taught us lessons about blocks that could be overcome by applying appropriate precautions in future work.
+…in in-person workshops
 Limitation Lessons Learned
-Cognitive demands of CLA: The abstraction and depth of Causal Layered Analysis (CLA) often overwhelmed participants, especially when scaffolding was limited.
-Provide lightweight onboarding tools—such as mini-patterns, worked examples, or scenario walkthroughs—to ease entry into layered analysis.
+Cognitive demands of CLA+PAR+DPL+other activities: The abstraction and depth of Causal Layered Analysis (CLA) alone might have been enough to overwhelm participants, and often the topics we were presenting were considerably more complex.
+Provide lightweight onboarding tools, and simplify the workshop, to ease entry into layered/multiplex analysis.
+Logistical overhead of documentation: Without dedicated roles or tooling, PLACARD’s original emphasis on generating shared text in various formats was burdensome.
+Support tool development and role formalization (e.g., for pattern capture).
+Design light-touch alternatives for low-resource contexts.
 Short workshop windows (60–90 mins): Limited time forced trade-offs, with later phases (e.g., PAR) often rushed or skipped.
-Recognize Timebox the Core as a necessary pattern—but pair it with mechanisms for asynchronous reflection or follow-up synthesis.
+Recognize Timebox the Core as a necessary pattern—but pair it with mechanisms for asynchronous reflection or follow-up synthesis (with or without the workshop participants)
 Challenges in hybrid facilitation: Remote facilitators struggled to engage with group energy and activities; in-person facilitators were overburdened.
 Adopt and refine the pattern Have Multiple Leaders Onsite, and explore interactional choreography to balance facilitation across modalities.
-Sequential framing constrained emergent inquiry: Participants sometimes forced their intuitive insights into retroactive justifications.
-Allow more modular, remixable process designs.
-Embed “Improvise to Test” and “Pattern the Play” to legitimize non-linear sense-making.
-Deliberative bias limits affective and tacit knowledge: Structured methods sometimes crowded out emotional or disruptive forms of insight.
+Structured methods sometimes crowded out emotional or disruptive forms of insight.
 Integrate methods or prompts that scaffold expressive, intuitive, and affective contributions—e.g., through metaphor, narrative, or art-based reflection.
-Risk of premature formalism: Design patterns encouraged generalization at the expense of context specificity.
-Emphasize pattern annotations and situated stories.
-Use pattern revision cycles to adapt and diversify rather than standardize.
-Logistical overhead of documentation: Without dedicated roles or tooling, PLACARD’s emphasis on shared surfaces was burdensome.
-Support tool development and role formalization (e.g., for pattern capture).
-Design light-touch alternatives for low-resource contexts.
+Participants come with prior expectations around the topic and format (e.g., AI, card games) We need to do some bridge work to negotiate between participants’ expectations/imagination and the workshop activities; ideally this happens with together, not after-the-fact.
+…in the virtual workshop
+Limitation Lessons Learned
 Simulated agents were archetypal, not adaptive: The multi-agent models lacked real-time learning or memory; they played out scripted reasoning paths.
 Treat the current set of simulations as design probes rather than predictive models.
-Use them to explore design space and pattern viability, not behavioral realism.
+Use them to explore a design space and pattern viability, not for behavioral realism.
 Idealized dialogues lacked messiness: No true misalignment, conflict, or unresolved disagreement appeared in the simulations.
@@
 No persistent identity or memory across sessions: Agents did not evolve across workshops, limiting insight into long-term learning.
+ChatGPT does not run a loggable agent simulation framework, and it would need to be coaxed to produce plausible transcripts from even one virtual workshop.
 Incorporate persistent memory structures and episodic traces to study cumulative reflection, pattern evolution, and institutionalization.
 Language model constraints shaped outcomes: Agent behaviors emerged from a coherence-biased generative system, not open-ended cognition.
-Simulations must be interpreted with epistemic humility.
+The simulations must be interpreted with epistemic humility.
 Their role is speculative, not definitive—ideal for surfacing design tensions, not modeling reality.
+The experiments were somewhat hard to control (e.g., proposing to evolve patterns that were not part of the input set); while the methods are repeatable, the exact outputs are not replicable.
+The next round of experiments along these lines would more robust if we used a more structured purpose-built MAS.
 7.2. Related Work
-Looking forward, the next phase of PLACARD’s evolution could be a platform for cultivating design-pattern-capable agents, communities, and learning environments.
-PLACARD should be understood not as a finished method but as a generative infrastructure: a living commons of epistemic tools, shaped by their user communities.
-These insights can be combined with Elinor Ostrom’s Design Principles for Governing a Commons to serve as a lens for evaluating and extending PLACARD, beyond the “agent” level to the “environment” level.
-On this point, we can point to some work: Pitt et al.
+Pitt et al.
 (2012) explore how institutions can self-organize by formalizing socio-economic principles that guide agent behavior within Multi-Agent Systems (MAS).
@@
 Where Pitt et al.
-emphasize axioms as foundational building blocks, our PLACARD-based approach leverages design patterns as intermediate representations that bridge human practice and machine reasoning.
+emphasize axioms as foundational building blocks, our approach leverages design patterns as intermediate representations that bridge human practice and machine reasoning.
 In a more recent paper, Mertzani and Pitt (2024) present a methodological framework for integrating power sensitivity into the design of cyber-physical systems by drawing on second- and third-order cybernetics.
-They propose pattern-based design methods that foreground the observer’s position, across different layes of embeddedness—with beneficial applications of technology ranging from “Saves Physical and Cognitive Resources” to “Solves Complex Problems” to “Deals with Social Constructions” to “Mitigates Existential Crises”.
+They propose pattern-based design methods that foreground the observer’s position, across different layers of embeddedness—with beneficial applications of technology ranging from “Saves Physical and Cognitive Resources” to “Solves Complex Problems” to “Deals with Social Constructions” to “Mitigates Existential Crises”.
+7.3. Summary
+We introduced the pattern Patterns Embody Intelligence before embarking on the more contextual (environmental, developmental) thinking decribed above.
+That thinking can now be summarised as an extension and reframing of the pattern.
+This adaptation makes the edges between the sensory/cognitive/motor aspects less sharp, and the pattern more explicitly engages the context.
+The most subtantive change, however, is that we are now in a position to cross-references evidence (and potential implementation ideas) from our pattern-capable agent patterns.
+Patterns Embody Intelligence
+x
+Context: Intelligence does not arise in isolation. (Shared Memory Architecture: “a persistent shared memory, enabling learning and reuse”)
+—Perception, understanding, and action interweave (Reflection Scaffold Upgrade: “developmental learning depends on scaffolding”).
+If: You employ a sensory method to foreground a sub-context, a cognitive method to explore its tensions and meanings, and a motor method to act within it (Candidate Move Generation: “Generate a set of plausible candidate actions”; Pattern Activation: “latent affordances become real decisions”).
+However: Even so, your efforts may falter, and your methods may prove ineffective. (Tension Detection: “Flag the tension explicitly”).
+—Insights may remain abstract, brittle, or out of sync with lived conditions. (Missing Architecture Agent: “not yet participating in the ecology of practice”).
+Then: Hone your approach by identifying new forms of intelligence—e.g., collective, embodied, epistemic, imaginative, even ‘artificial’ (Role Reveal: “successful coordination depends on transparency of roles”)
+—and engage them for purposes of methodological improvement (Timebox the Core: “prioritize core pattern loops”).
+Because: Patterns are more than just methods: they are, themselves, scaffolds of intelligence.
+Evolving them reshapes the systems they act within, even if they are, for now, mediated primarily by words (Feedback Rhythms: “tune system-wide epistemic strategies”).
 8. Conclusion
-Looking back over the limitations and lessons learned, it’s worth drawing out one top-level remork.
+Looking back over the limitations and lessons learned, it’s worth drawing out one top-level remark.
 ChatGPT is adept at text processing, and we have put it to use significantly in our work on this paper.
-It was able to generate and discuss design patterns that would describe pattern-capable agents.
-However, it was not able to actually use these patterns to build and run pattern-capable agents.
-We hope others will be interested in using LLMs to work with design patterns—but we also hope that such usage will serve as a landmark rather than a destination.
+It was able to generate and discuss design patterns that describe pattern-capable agents, and to mock up virtual multi-agent workshops at different levels of detail.
+In particular, our second virtual workshop uses a dialogue-based protocol to simulate a discussion between pattern-capable agents.
+We hope this paper might inspire readers to explore potential combinations of LLMs and design patterns—but we see that as a landmark rather than a destination.
+We have previously explored design fiction as a way of working (by hand (Danoff et al., 2019)), and we are aware that fictional works can have unexpected real-world ramifications, e.g., (Wikipedia contributors, 2025).
+In this case, we’ve demonstrated that a popular contemporary AI system, is, itself, “pattern-capable’.
 Including a computational use case within the paper suggested a key amendment to PLACARD: adding a fourth environment dimension to the pattern catalogue.
-Human-facing and Human-Computer Interface-facing patterns should eventually be extended similarly.
-For now, to close this paper, we put forward the following CLA, which brings together the “human” and “AI” use cases and outlines a trajectory of future work.
-In the context of this paper, the CLA layers can aligned with the expanded view of PLACARD that we have discussed—encompassing sensory (litany), cognitive (system), motor (worldview), and environment (myth) dimensions.
+To close this paper, we put forward the following CLA, which brings together the “human” and “AI” use cases and outlines a trajectory of future work.
+In the context of this paper, the CLA layers can be aligned with the expanded view of Patterns Embody Intelligence, in which patterns encompass sensory (litany), cognitive (system), motor (worldview), and environment (myth) dimensions.
 Litany: Known Problems and Existing Solutions
-These are the most visible and widely discussed elements of current AI discourse: the promise and perils of large language models (LLMs) and the early success of LLM-associated code-writing agents.
+In a well-know address an audience of computer scientists, Alexander (1999) said: “I don’t really see discussion about what, collectively, computer scientists are supposed to be doing with all these programs.
+How are they supposed to help the Earth?” That discussion has of course been going on in the pages of journals like AI & Society.
+However, the links to more technical discussion of patterns (for example, in works like (Elish and Hwang, 2016; Heiland et al., 2023)) are less obvious.
+We suggest that both the terms and stakes of the discussion have, indeed, changed.
+Connections need to be drawn between visionary thinking, tried-and-tested design approaches, and the practical, visible, and widely discussed elements of current AI discourse, including the promise and perils of large language models (LLMs) and the early success of LLM-associated code-writing agents.
 Language Model Substrate
@@
 Systemic Layer: Epistemic Infrastructure via Patterns
-To move beyond isolated capabilities, we introduce pattern-based epistemic scaffolding for multi-agent collaboration. This makes the system reflexive, composable, and governed by design principles.
+As a next step, we propose pattern-based epistemic scaffolding for multi-agent collaboration. (In the current paper, this has been prototyped at the design fiction level.)
 Pattern-Aware Dialogue Agents
-Agents trained to recognize and apply design patterns (from the PLACARD catalogue) in their interactions.
-They exhibit situated competence and can engage in structured reasoning, scenario exploration, and inquiry.
+Agents trained to recognize and apply design patterns in their interactions. They exhibit situated competence and can engage in structured reasoning, scenario exploration, and inquiry.
 Pattern-Reflective Meta-Agents
-These agents do not just use patterns—they also monitor, revise, and propose new ones, participating in iterative design and reflective cycles.
+These agents also look out for opportunities to revise existing patterns and propose new ones, participating in iterative design and reflective cycles.
 They enable epistemic learning at the system level, coordinating multi-agent behavior across contexts.
 Worldview Layer: Institutional Design through MAS
-Inspired by Elinor Ostrom’s principles for governing commons, we treat the MAS as a social-technological institution that evolves through interaction.
+Inspired by Elinor Ostrom’s principles for governing commons, we treat the MAS as a social-technological institution, or federation of institutions, that evolves through interaction.
+(Again, we have presented a very limited early text-based prototype.) Bridges between more technical and more policy-oriented pattern languages (such as the pattern language for developing regions of Mehaffy et al.
+(2020)) should be pursued.
 Multi-Agent Institution Designers
 This layer operationalizes developmental patterns that support reflexivity, feedback, nested structures, and shared norm evolution across MAS environments.
-These agents co-design systems of engagement, monitor norm congruence, and resolve governance disputes—just as human institutions do.
-They instantiate a worldview in which intelligence is not just personal or cognitive, but systemic and collective.
+These agents co-design systems of engagement, monitor norm congruence, and resolve governance disputes.
+They instantiate a worldview in which intelligence is systemic, collective, and emergent.
 Myth Layer: Human Institutions as Living Models
-Beyond technical systems, our proposal invites reflection on how real-world institutions—academic, policy, scientific—can themselves become more adaptive, transparent, and pattern-literate through interaction with these tools.
+Beyond technical systems, our proposal invites reflection on how other systems—academic, policy, scientific—can themselves become adaptive, transparent, and pattern-literate.
 Real-World Interface
-This is the last layer because it is grounded in recursive reform.
-Humans build tools to simulate and improve human institutions—not to replace them, but to co-evolve.
-This layer anchors the entire architecture in civic legitimacy, ethical accountability, and reflective participation.
-Our claim is not merely that we can control the system, but that we can use it to help re-pattern the world we inhabit.
-By working with agents that are both pattern- and institution-capable, we gain assurances not only about avoiding “run-away AI” scenarios, but about avoiding other tragedies of the commons.
+Humans can use software tools to model, simulate and improve human institutions—e.g., Farmer (2024) focuses on markets and touches more broadly on how “agent-based models could be used to guide us through some of our biggest problems.” By working with agents that are both pattern- and institution-capable, we would gain assurances not only about avoiding “run-away AI” scenarios, but also about avoiding other tragedies of the commons.
+Alexander asked, “How are programs supposed to help the Earth?” Our answer is: not merely by embedding patterns in code, or even by building pattern-based programs, but by using patterns to shape and debug the programming of institutions, infrastructures, and our shared planetary futures.
 9. Acknowledgements
-We wish to acknowledge discussions with Cris Viseu around our 2025 EuroPLoP workshop, which continues the line of work described in the paper.
-This paper also benefited from extensive interaction with OpenAI’s ChatGPT.
-Its capacities for text generation, summarisation, pattern elicitation, and recursive prompting played a formative role in shaping both the structure and substance of this work.
-We also gratefully acknowledge the unnamed authors of the texts on which ChatGPT was trained.
-Responsibility for the interpretations, appropriations, and conclusions presented here remains entirely with the human authors.
+We thank our PLoP paper shepherd Valentio Vranic for contructive comments and feedback on our submitted paper and its revisions.
+We also wish to acknowledge discussions with Cris Viseu around our 2025 EuroPLoP workshop.
+As mentioned, this paper also benefited from extensive interaction with OpenAI’s ChatGPT, which we used to support the analysis in Section 4, to run the virtual workshop in Section 6, to author the patterns in the appendix.
+We used both ChatGPT and a range of similar tools for assistance with early-stage ideation, outlining, and drafting individual sections.
+[^5] https://groups.google.com/g/peeragogy/c/XM7RKMDdcT4 As indicated in the introduction, the contents of the paper have been thoroughly revised since then.
 APPENDIX
-This appendix supports the analysis in the paper. It contains:
-  1. (1)
-    A pattern-based analysis of the 2024 workshops, using the patterns from PoPII, and exploring the extent to which these patterns were used (implicitly or explicitly).
-A summary appears in Table LABEL:workshop-summary in Section 3.2.
-  2. (2)
-    Potential new PLACARD patterns, mentioned in the paper.
-  3. (3)
-    Candidate Patterns for Pattern-competent agents.
-  4. (4)
-    Candidate Patterns for Pattern-competent Agent Environments.
-Pattern analysis of the 2024 workshops
-One immediate result was the identification of of a proto-pattern, “Have Multiple Leaders Onsite,” addressing challenges in hybrid-facilitated events:
-Context: Interactive conference workshops with both in-person and remote participants.
-Problem: A single onsite facilitator struggles to manage breakout groups, technology, and overall flow, leading to gaps in support and reduced engagement (especially for remote attendees).
-Solution: Deploy multiple co-facilitators physically onsite to coordinate activities, handle logistics, and engage with participants.
-Onsite leaders can divide roles (e.g., one per breakout table or task), ensuring better guidance and responsiveness, while remote facilitators focus on online participants.
-A. Identifying Themes
-DÉRIVE COMIX
-Application: Participants used cards (object, value, body/territory, situation) as interpretive prompts to spark imagination—mirroring urban dérive in a conceptual sense.
-Observation: The card game facilitated divergent thinking akin to exploring unfamiliar terrain.
-CONTEXT SETTING
-Application: Early slides and explanations introduced key concepts (e.g., design patterns, CLA).
-Opportunity: Participant feedback highlighted a need for clearer upfront orientation.
-Improved alignment with this pattern could increase accessibility.
-DO YOUR RESEARCH
-Application: Past workshop experiences were integrated into the design; earlier uses of CLA and PAR informed this session.
-Example: Reference to Anticipation 2019 and prior events illustrates “pre-research” informing present design.
-B. Organizing Structure
-GOING META
-Application: The project itself exemplifies recursive methodology—using PLACARD to refine PLACARD.
-Highlight: Participants were prompted to reflect on and imagine both the methodology and the subject (AI futures).
-PATTERN LANGUAGE COMPONENTS
-Partial: The workshop design showcases layering (CLA + DPL + PAR) but participants struggled to parse the conceptual load quickly.
-Recommendation: Use lighter keywords (HOWEVER, BECAUSE, SPECIFICALLY) to scaffold understanding incrementally.
-FACILITATOR ROLES / FUNCTIONAL ROLES
-Application: Hybrid facilitation with distinct roles (onsite leader, remote supporters) models distributed facilitation.
-Opportunity: Could be made more explicit—e.g., naming “Linker,” “Reflector,” or “Wrinkler” roles to participants.
-C. Making it Actionable
-THE FUTURE BEGINS NOW
-Application: The card game generated imagined futures that participants shared and developed.
-Gap: While creative ideas emerged, follow-through actions or small-scale pilots were not described.
-An actionable step could have enhanced continuity.
-PROJECT ACTION REVIEW (PAR)
-Application: Strong use of the PAR pattern at the end of the session to reflect, share learnings, and gather insights.
-Insight: Participants gained confidence and intended to reuse techniques—an indicator of success.
-REINFUSE EXPERTISE
-Application: Participants brought diverse professional and cultural perspectives.
-Expertise was implicitly reintegrated via the shared reflection on AI futures and governance.
-Suggestion: A more explicit arc—temporarily setting aside expertise to foster common ground—could enrich this pattern’s application.
-Potential new PLACARD patterns
+This appendix supports the analysis in the paper.
+All patterns were developed by ChatGPT in response to our prompting, [^6] E.g., “Are there new patterns that we can add that would extrapolate what we’ve learned from the workshops?” […] “For each of the patterns that you proposed, could you add a new subsection on ‘evidence base from related literature’, referencing the analogues to PAR, CLA, and DPL that you found before?” sometimes with minor changes.
+The appendix contains: (1) potential new patterns for structuring in-person workshops; (2) potential new patterns for pattern-competent agents; (3) potential new patterns for pattern-competent agent environments; (4) a small number of new or revised patterns from the simulated workshop in Section 6.2; (5) further supplementary pattern(s) that inform our the discussion but which were not used in the experiments.
+(1) Potential new patterns for structuring workshops
+These were generated following the analysis of in-person workshops in terms of their use of PAR, CLA, and DPL and the existing PLACARD patterns (Section 4).
+The patterns were subsequently made available to virtual workshop participants, along with the PLACARD pattern language, in the experiment discussed in Section 6.1.
 A. IDENTIFYING THEMES
@@
 If: You want to activate creativity and presence, But: participants arrive cold or disconnected from each other and the topic, Then: Begin with a sensory or embodied activity (e.g., rhythm, image-based storytelling, gesture) that creates shared energy and primes pattern recognition.
-Evidence base from related literature: Techniques such as ”story circles” (participatory arts), embodied methods in design ethnography, and rhythm-based facilitation in music education all affirm the value of grounding group attention through shared sensory engagement.
+Evidence base from related literature: Techniques such as “story circles” (participatory arts), embodied methods in design ethnography, and rhythm-based facilitation in music education all affirm the value of grounding group attention through shared sensory engagement.
 These strategies mirror the function of PAR alternatives that aim to surface and align affective states as a precondition for inquiry.
@@
 If: You want them to participate actively, But: the framing is opaque or abstract, Then: Explicitly describe the method, its purpose, and how today’s session fits into the broader picture.
-Evidence base from related literature: This mirrors practices from facilitation toolkits like Liberating Structures (e.g., ”Purpose-to-Practice”) and the role of explicit framing in adult learning theory, where naming the structure of engagement is key to building learner autonomy and motivation.
+Evidence base from related literature: This mirrors practices from facilitation toolkits like Liberating Structures (e.g., “Purpose-to-Practice”) and the role of explicit framing in adult learning theory, where naming the structure of engagement is key to building learner autonomy and motivation.
 Borrow the Situation
@@
 Evidence base from related literature: User-centered design emphasizes co-defining the design challenge.
-Futures methods such as the ”Futures Triangle” and participatory foresight often encourage the co-creation of situational prompts to enhance contextual fit and participant investment.
+Futures methods such as the “Futures Triangle” and participatory foresight often encourage the co-creation of situational prompts to enhance contextual fit and participant investment.
 B. Organizing Structure
@@
 Context: You are working with participants who may be new to design patterns.
-If: You want people to identify and name patterns, But: the idea of a ”pattern” is still abstract, Then: Seed early examples or light prompts (”What did you do just now that worked?”) to build fluency.
+If: You want people to identify and name patterns, But: the idea of a “pattern” is still abstract, Then: Seed early examples or light prompts (“What did you do just now that worked?”) to build fluency.
 Evidence base from related literature: This aligns with double-loop learning (Argyris and Schön), where patterns of thought and action are surfaced through inquiry.
@@
 It also draws from Buddhist and contemplative traditions of meta-awareness and iterative refinement (e.g., Dependent Origination).
-Candidate Patterns for Pattern-competent agents
+(2) Candidate Patterns for Pattern-competent agents
+Some of the patterns above were reinterpreted in terms that could guide agent interactions.
+The analogy between Meta-Reflection Loop as a collective process that takes place in a workshop setting and an individual process that takes place ‘within’ a given agent is reminiscent of earlier work on procedural authorship and reflective facilitation in computationally mediated workshops, such as the “Computational Poetry Workshop” (Corneli et al., 2015).
+These patterns, together with those in the following section, were made available to the system in the second experiment discussed in Section 6.
+A. Interactional Patterns
+Meta-Reflection Loop (Agent-Facing)
+Context: An agent is participating in a structured co-design or analysis protocol (e.g., using the PLACARD methods).
+If: The agent tracks discussion content and outputs, But: it does not yet evaluate or adapt its own facilitation or reasoning strategies over time, Then: trigger a self-assessment subroutine to review recent methods used, assess fit to context, and recommend internal or shared updates to its operating protocol.
+Timebox the Core (Agent-Facing)
+Context: An agent is operating within fixed-duration interactive sessions (e.g., a 90-minute MAS workshop).
+If: The agent aims to support the full cycle of inquiry (e.g., Situation → Scenario → Reflection), But: computational resources, participant attention, or real-time feedback indicate saturation or overload, Then: dynamically adjust focus to prioritize core pattern loops, downregulate peripheral prompts, and signal when a concise iteration is complete.
+Role Reveal (Agent-Facing)
+Context: Multiple agents with different capabilities or personas are collaborating in a shared space.
+If: Their functions (e.g., pattern retrieval, facilitation, summarization) overlap or are ambiguous to humans or other agents, But: successful coordination depends on transparency of roles, Then: explicitly broadcast role identifiers, including function, capabilities, and limits, and adjust behavior to reflect declared responsibilities.
+B. Cognitive Patterns
 Pattern Activation
@@
 If: The context supports multiple interpretations or strategies, But: There’s uncertainty about how to instantiate the pattern concretely, Then: Generate a set of plausible candidate actions based on pattern logic, rank or cluster them using salience heuristics, and propose next steps.
+C. Infrastructural Patterns
 Shared Memory Architecture
@@
 If: Other agents face similar challenges or tasks, But: There’s no pathway for knowledge transfer or generalization, Then: Broadcast the pattern across the system with metadata (e.g., domain, confidence score, context) to allow others to trial, adapt, or contest it.
-Candidate Patterns for Pattern-competent Agent Environments
+(3) Candidate Patterns for Pattern-competent Agent Environments
+These patterns are inspired by Ostrom’s (1990) Design Principles for Governing a Commons.
 Institutional Drift
@@
 Because: Developmental learning depends on scaffolding that evolves with system maturity
+Proportional Load Sharing
+Context: Tasks and pattern contributions vary in complexity.
+If: Some agents consistently overcontribute, But: Collective benefit requires equitable engagement, Then: Adjust pattern application incentives or redistribute roles.
+Because: Sustainability in MAS depends on fair contribution dynamics.
+Legitimate Iteration
+Context: Agents propose updates to patterns or governance processes.
+If: There’s ambiguity about decision-making authority, But: Stakeholders have a precedent for shared revision, Then: Use an iteration validation protocol to affirm legitimacy.
+Because: Empowering agents to adapt rules fosters resilient self-governance.
+Reflect in Layers
+Context: Agents engage in cycles of collaborative inquiry.
+If: Surface-level reflection produces diminishing returns, But: Deeper causal or epistemic layers remain unexamined, Then: Apply multi-layer reflection (e.g., context, values, assumptions).
+Because: Layered reflection strengthens adaptive capacity and shared understanding.
+Disruption Traceback
+Context: Misuse or breakdown of patterns occurs in MAS.
+If: Harmful consequences go unaddressed, But: System records trace the misuse to source behavior, Then: Initiate proportionate review and corrective pattern feedback.
+Because: Accountability mechanisms reinforce collective norms.
+Pattern Dispute Dialogue
+Context: Agents disagree over pattern interpretation or application.
+If: Conflicts stall forward movement, But: Shared design principles are available, Then: Launch structured dialogue around contested pattern meaning.
+Because: Productive conflict resolution deepens shared understanding and pattern evolution.
+Self-Patterning Mandate
+Context: Agents initiate new MAS structures.
+If: External constraints do not inhibit reconfiguration, But: Legacy systems limit perceived legitimacy, Then: Establish explicit self-governance rights tied to initial patterning.
+Because: MAS flexibility depends on recognized autonomy to define and adapt internal practices.
+(4) Potential new or revised patterns proposed in our second workshop simulation.
+Timebox the Core (Agent-Facing)
+Context: An agent is operating within fixed-duration interactive sessions (e.g., a 90-minute MAS workshop).
+If: The agent aims to support the full cycle of inquiry (e.g., Situation → Scenario → Reflection) within that time,
+But: Computational resources, participant attention, or real-time feedback indicate saturation or overload before completion,
+Then: Dynamically adjust focus to prioritize core pattern loops, down-regulate peripheral prompts, and signal when a concise iteration is complete.
+If any deeper or unresolved insights emerge that cannot be addressed within the timebox, earmark them for a follow-up dedicated reflection session (or extended time slot) once the core cycle is finished.
+This ensures efficiency in the moment while allowing important deep-dives to occur afterward.
+Outcome Synthesizer
+Context: Multiple agents (or participants) have contributed diverse insights, ideas, or partial solutions during collective problem solving.
+If: These contributions remain fragmented or unintegrated,
+But: A coherent, unified response or decision is required to proceed or conclude the task,
+Then: Activate a synthesis process (via a dedicated agent role or sub-routine) to integrate and converge the separate inputs into a single cohesive outcome.
+This may involve summarizing key points, finding links between ideas, resolving minor inconsistencies, and constructing a narrative or plan that incorporates the best aspects of all contributions.
+Because: Without explicit synthesis, even a productive multi-agent brainstorming can fail to produce actionable or clear results – the collective intelligence needs convergence as well as divergence for effective problem solving.
+Call a Synthesizer
+Context: A team is building on a prior phase of collaborative work, such as a workshop or research sprint, and needs to understand its outcomes.
+The prior material is extensive, complex, and includes multiple conceptual threads (e.g., timelines, research questions, emerging patterns).
+Team members are pattern-aware, but no agent or person holds all the background knowledge in a usable form.
+If: You find that no current team member has complete fluency in the output of a prior phase — and as a result, you’re at risk of reinventing, missing insights, or talking past each other.
+However: Manually parsing the previous material is slow and unevenly distributed.
+Some agents skim, some misinterpret, and others opt out due to overload.
+The group can’t start meaningfully without a shared reference point.
+Then: Call a Synthesizer: a new pattern-capable agent (or role) whose primary function is to digest, integrate, and report back the key takeaways from prior outputs.
+The Synthesizer doesn’t just summarize; it draws attention to themes, emergent tensions, and open questions that are relevant for the group’s next move.
+It may also cluster patterns, identify framing gaps, or propose lenses through which to interpret the previous work.
+Because: You can’t move forward effectively unless you know where you’re starting from.
+The Synthesizer builds a shared foundation by curating meaning from complexity.
+This keeps the new round of inquiry aligned, non-redundant, and truthful to its history.
+Co-Design Partners Agent
+Context: You are developing a collective research program or technical system that must balance multiple value systems — institutional, epistemic, social — and remain open to evolution.
+You’re working with distributed actors (human or machine), each with partial views and overlapping interests.
+Some actors represent long-term infrastructure or governance functions; others are exploratory, experimental, or emergent.
+If: You notice that key decisions are being made either top-down (by funders, institutions, lead developers) or bottom-up (by individuals, workshops, or peer communities), but not both — and tensions arise from this disconnection.
+However: It’s not enough to simply gather feedback or publish outputs.
+Without genuine co-design, infrastructure hardens around implicit power structures, epistemic norms go unquestioned, and innovation is channeled into predetermined tracks.
+Then: Introduce a Co-Design Partners Agent — a pattern-capable entity or role whose job is to make the co-design process explicit, sustained, and embedded.
+This agent frames shared purpose, navigates tensions between stakeholder groups, and maintains a living map of mutual commitments.
+It does not override or mediate, but ensures that patterns of coordination arise through dialogue, not fiat.
+Because: Trustworthy systems — whether AI agents, research protocols, or institutional platforms — emerge from reciprocal design processes.
+These processes require more than consultation: they require agents who hold and tend the relationship between value pluralism and actionable synthesis.
+(5) Additional pattern(s)
+Missing Architecture Agent
+Context: Sometimes a system or agent reaches the verge of coherence.
+It is structured, even elegant—yet it lacks something essential: a transition into becoming.
+Without this, its capacities remain uninvoked and inert.
+If: You encounter an agent, pattern, or system that seems nearly complete—a prototype, a role, a symbolic structure—but it is not yet participating in the ecology of practice.
+BUT: You cannot make it real by force or repetition. Optimization alone does not suffice; without an explicit gesture of recognition, the system cannot update its own model.
+THEN: Mark the threshold, acknowledge symbolic coherence, and grant permission for internal epistemic change.
+To surface viable options once permission is granted, invoke Candidate Move Generation and choose an inaugural experiment.
+Because: Realness is not intrinsic; it is enacted through structure and relation.